### v3

Abstract.
Empirical evidence shows that wealthy households have substantially higher saving rates and markedly lower marginal propensity to consume (MPC) than other groups. Existing theory cannot account for this pattern unless under restrictive assumptions on returns, discounting, and preferences. This paper develops a general theory of optimal savings with preference shocks, allowing risk aversion to vary across states and over time, and shows that incorporating such heterogeneity in risk attitudes fundamentally reshapes the asymptotic dynamics of consumption and saving. In particular, zero asymptotic MPCs (100% asymptotic saving rates) arise under markedly weaker conditions than in existing theory. Strikingly, such outcomes occur whenever there is a positive probability that agents become less risk averse in the future. Therefore, the vanishing MPC emerges as a generic feature rather than a knife-edge result of the optimal savings model, offering a more theoretically robust and empirically consistent account of the saving behavior of wealthy households.
Keywords: MPC, saving rate, time-varying risk aversion, wealth inequality.
A Theory of Saving under Risk Preference Dynamics*\@makefntextWe thank Émilien Gouin-Bonenfant, Mi Luo, Nisha Peng, John Stachurski, Shenghao Zhu, as well as seminar audiences at CUEB and CUFE for helpful comments and suggestions. Qingyin Ma gratefully acknowledges financial support from NSFC (No.72573111).
Qingyin Ma\@makefntextISEM, Capital University of Economics and Business. Email: qingyin.ma@cueb.edu.cn.,  Xinxi Song\@makefntextISEM, Capital University of Economics and Business. Email: songxinxi@cueb.edu.cn.,  Alexis Akira Toda§\@makefntextDepartment of Economics, Emory University. Email: alexis.akira.toda@emory.edu.
1. Introduction
In recent decades, a large body of empirical evidence has shown that wealthy households save a substantially larger fraction of their income and exhibit markedly lower marginal propensity to consume (MPC) than other groups. For instance, Dynan et al. (2004) document a strong positive relationship between lifetime income and saving rates. Using Norwegian administrative data, Fagereng et al. (2025) find that households in the top 1% of the wealth distribution report saving rates far exceeding those of the median. Moreover, Mian et al. (2025) document the emergence of a saving glut of the rich, finding that since the 1980s, the top 1% of wealthy households in US has sharply increased its saving rate from about 43% to over 54% of disposable income, while saving rates for the rest of the population have declined.
Because household saving behavior directly drives wealth accumulation, and wealthy households hold a disproportionate share of total assets, [^1] Saez and Zucman (2016) document that the top 0.1% wealth share in the United States increased from 7% in 1979 to 22% in 2012. Moreover, Piketty et al. (2019) estimate that the top 10% wealth share in China rose from 40% in 1995 to 67% in 2015. their saving patterns have first-order implications for the long-run distribution of wealth and the mechanisms that sustain inequality (Benhabib and Bisin, 2018). From a macroeconomic perspective, even small variations in the MPC or risk-taking behavior of the wealthy can have outsized effects on aggregate saving rates and capital formation (Carroll et al., 2017; Lee, 2023), the equilibrium interest rates and asset prices (Mian et al., 2013, 2021; Patterson, 2023), the transmission of fiscal and financial shocks (Cho, 2023; Ampudia et al., 2024; Auclert et al., 2025), and the design of macroprudential policy (Jappelli and Pistaferri, 2014; Crawley and Kuchler, 2023).
Despite these empirical regularities, existing studies remain limited in explaining the persistent high saving rates and low MPCs of wealthy households. [^2] As is standard in the literature, we define the saving rate as the change in net worth relative to total income, excluding capital losses: $s_{t+1}=\frac{w_{t+1}-w_{t}}{\max\{(R_{t+1}-1)(w_{t}-c_{t}),0\}+Y_{t+1}},$ where $s_{t}$, $w_{t}$, $c_{t}$, $Y_{t}$, and $R_{t}$ denote the saving rate, wealth, consumption, nonfinancial income, and the gross rate of return on wealth, respectively. The asymptotic saving rate refers to the saving rate in the limit as $w_{t}\to\infty$. The canonical incomplete market model, as emphasized by Ma and Toda (2021), predicts that in a stationary equilibrium with constant relative risk aversion, discount factor, and return on wealth, the asymptotic saving rate is negative, contradicting the empirical evidence that the wealthy continue to accumulate assets indefinitely. Subsequent studies have introduced return risk and stochastic discounting to reconcile theory with data (Benhabib et al., 2015; Ma et al., 2020; Ma and Toda, 2021). However, even in richer stochastic environments, as Ma and Toda (2021) show, matching the observed low MPCs and high saving rates among wealthy households requires quantitatively stringent conditions, and the resulting implications remain fragile.
A key limitation of this literature lies in treating risk preferences as homogeneous and time-invariant. Recent research highlights that households differ not only in income, wealth, and investment opportunities, but also in their attitudes toward risk—and that these attitudes evolve over time. While the canonical savings model assumes a fixed coefficient of relative risk aversion, a growing body of empirical and experimental studies reveal that risk aversion is heterogeneous across individuals and varies systematically over the life cycle and across business cycles (Andersen et al., 2008; Cohen and Einav, 2007; Falk et al., 2018; Finkelstein et al., 2013). Moreover, fluctuations in risk tolerance comove with individual financial circumstances, macroeconomic conditions, and asset market volatility (Guiso et al., 2018; Malmendier and Nagel, 2011; Sahm, 2012; Schildberg-Hörisch, 2018), suggesting that agents continuously reassess their exposure to risk in response to both individual and aggregate shocks.
These observations motivate a reconsideration of the intertemporal saving behavior under dynamically evolving risk preferences. By abstracting from fluctuations in risk attitudes, existing studies overlook the endogenous interaction between risk perceptions, wealth accumulation, and saving behavior, potentially distorting predictions of long-run saving dynamics. [^3] From a theoretical perspective, the assumption of constant risk aversion plays a critical role in generating the negative asymptotic saving rates implied by standard models. With fixed preferences, precautionary motives weaken as wealth grows, causing consumption to eventually outpace income and leading saving rates to turn negative at high wealth levels. Allowing risk aversion to vary over time and across states, by contrast, can profoundly alter consumption and saving incentives, particularly for wealthy households, whose consumption and portfolio choices are especially sensitive to shifts in risk perceptions.
In this paper, we develop a general theory of optimal savings with preference shocks, allowing risk aversion to be both time-varying and state-dependent, and show that this generalization fundamentally reshapes the asymptotic dynamics of consumption and saving, offering a parsimonious and robust explanation for the persistently high saving rates and low MPCs observed among wealthy households.
To begin, we study a generic optimal savings problem, in which agents face preference shocks, and the rate of return, discount factor, and nonfinancial income are all allowed to be state-dependent, time-varying, and mutually dependent. We establish the existence and uniqueness of the optimal consumption policy in a suitable candidate space. Our key assumption admits a transparent interpretation in terms of the discounted average gross payoff on assets and is straightforward to verify in practice.
We then study the properties of the optimal consumption policy. Under general specifications, we establish its continuity and monotonicity, and analytically characterize the wealth threshold above which saving occurs. To explore the consumption and saving behavior of the wealthy households, we adopt a generalized CRRA framework. Departing from the standard formulation, we allow the level of risk aversion to vary with the state. This state-dependent specification aligns with the empirical evidence that individuals adjust their risk tolerance in response to changes in preferences, financial circumstances, and prevailing economic conditions, providing a more realistic foundation for studying intertemporal consumption and saving behavior.
We show that incorporating such preference heterogeneity fundamentally changes the asymptotic dynamics of saving. In particular, we provide an analytical characterization of the asymptotic MPCs and demonstrate that zero asymptotic MPCs, which correspond to a 100% asymptotic saving rate, emerge under markedly weaker conditions than in the benchmark model without preference shocks. Strikingly, whenever agents face a positive probability of transitioning to a lower risk aversion state, the precautionary saving motive remains perpetually active, regardless of how much wealth has been accumulated. [^4] Indeed, the prospect of becoming less risk averse in the future induces additional precautionary saving today, as agents smooth consumption across potential preference states. In this case, consumption grows more slowly than wealth, and the asymptotic MPC converges to zero. Importantly, this outcome is entirely independent of stochastic discounting, return risk, or income risk, and thus represents a highly general property of asymptotic saving behavior.
The asymptotic MPC is strictly positive at some state only if the probability of transitioning to a lower risk aversion state from that state is zero. In this case, risk attitudes evolve one-sidedly—risk aversion may remain constant or increase but cannot decline—and positive asymptotic MPCs may arise, depending on the interaction between the stochastic discount factor and the return on wealth. [^5] For the asymptotic MPCs to remain positive across all states, however, the transition matrix of risk aversion must be upper triangular, implying that agents can only become weakly more risk averse over time, which is empirically implausible. In the extreme case where agents always transition to a higher risk aversion state (from some state), the asymptotic MPC converges to one (at that state), and the consumption function becomes nonconcave in wealth, reflecting an extreme form of precautionary behavior.
We then complement our theoretical analysis with a quantitative investigation. First, we provide numerical illustrations that render the asymptotic results transparent and economically intuitive. Second, following Koop et al. (1996), we compute state-dependent impulse response functions to uncover the economic mechanisms through which preference shocks reshape consumption and saving dynamics.
Our quantitative analysis confirms that shocks to risk aversion generate economically significant and persistent responses in consumption and saving. Importantly, both the magnitude and persistence of these responses depend on households’ current state and on the persistence of preference shocks. In particular, the effects vary systematically with wealth and the prevailing level of risk aversion, revealing pronounced nonlinearities in the transmission of preference shocks across the wealth distribution.
Taken together, our framework provides a theoretically robust and empirically consistent account of the saving behavior of wealthy households, highlighting how risk preference heterogeneity offers a natural mechanism for the persistence of wealth accumulation. While we do not explicitly model wealth inequality, our analysis points to a novel mechanism that may be relevant for understanding wealth distribution in richer environments.
Related Literature
Early studies of the asymptotic saving behavior trace back to the literature on the concavity of consumption functions. Carroll and Kimball (1996) show that, under hyperbolic absolute risk aversion (HARA) preferences, the consumption function is concave in wealth, implying asymptotic linearity of consumption with respect to wealth. Carroll (2004) further emphasizes this asymptotic linearity and note that zero asymptotic MPCs can arise in a standard optimal savings model when the interest rate is negative.
Building on these insights, Benhabib et al. (2015) characterize the asymptotic MPCs analytically in an optimal savings model with capital income risk, where the return to wealth follows an iid process. Ma and Toda (2021) generalize this framework by allowing both the rate of return and the discount factor to be state-dependent and jointly driven by a generic Markov process. Their contribution lies in delivering a complete characterization of the asymptotic MPC in this setting and identifying conditions under which optimal consumption is asymptotically linear in wealth. In their model, zero asymptotic MPCs arise only as a knife-edge case, contingent on restrictive assumptions. Their empirical analysis indicates that achieving a zero asymptotic MPC requires stringent assumptions on return risk, stochastic discounting, and key parameters such as risk aversion, limiting the empirical plausibility of their model.
We extend this literature by incorporating preference shocks into a generic optimal savings framework, allowing for time-varying and state-dependent risk aversion. The models in previous studies, such as Ma and Toda (2021), emerge as special cases. In sharp contrast to these studies, the vanishing asymptotic MPCs are a generic feature of our framework rather than a knife-edge case. Furthermore, when risk aversion can only drift upward, consumption becomes nonconcave in wealth and the asymptotic MPC converges to one, departing sharply from the classical concavity-based intuition of Carroll and Kimball (1996), Ma et al. (2020), and Ma and Toda (2021).
Our study also builds on the theoretical literature on the optimal savings problem, a cornerstone of modern macroeconomics (Chamberlain and Wilson, 2000; Kuhn, 2013; Li and Stachurski, 2014; Benhabib et al., 2015; Ma et al., 2020). We extend the method of Li and Stachurski (2014) and Ma et al. (2020) to solve a generic savings problem with preference shocks. While this literature focuses primarily on existence, uniqueness, and computability, our analysis centers on the asymptotic properties of consumption and saving, which has no direct counterpart in these studies.
The rest of the paper is organized as follows. Section 2 formulates a generic optimal savings problem with preference shocks, and establishes existence, uniqueness, and key properties of the optimal policy. Section 3 systematically studies its asymptotic properties and discusses the resulting implications for saving and wealth accumulation. Section 4 presents a quantitative analysis that provides further intuition for our theory. Proofs and detailed discussion of the related algorithms are relegated to the Appendix.
2. Optimality results
To capture preference heterogeneity and its evolution over time, we study a general optimal savings problem with preference shocks.
2.1. Problem Statement
Agents choose a consumption-asset path $\{(c_{t},w_{t})\}$ to
(1) maximize $\mathbbm{E}_{0}\left\{\sum_{t=0}^{\infty}\left(\prod_{i=0}^{t}\beta_{i}\right)u(c_{t},Z_{t})\right\}$
(2) subject to $w_{t+1}=R_{t+1}(w_{t}-c_{t})+Y_{t+1}\quad\text{and}\quad 0\leq c_{t}\leq w_{t}.$
Here, the initial condition $(w_{0},Z_{0})=(w,z)$ is given, $u$ denotes the utility function, $\{Z_{t}\}_{t\geq 0}$ is a time-homogeneous Markov chain taking values in a finite set $\mathsf{Z}$, $\{\beta_{t}\}_{t\geq 0}$ is the discount factor process with $\beta_{0}=1$, $\{R_{t}\}_{t\geq 1}$ is the gross rate of return on wealth, and $\{Y_{t}\}_{t\geq 1}$ is non-financial income. These stochastic processes satisfy
(3) $\beta_{t}=\beta\left(Z_{t-1},Z_{t},\varepsilon_{t}\right),\quad R_{t}=R\left(Z_{t-1},Z_{t},\varepsilon_{t}\right)\quad\text{and}\quad Y_{t}=Y\left(Z_{t-1},Z_{t},\varepsilon_{t}\right),$
where $\beta$, $R$, and $Y$ are nonnegative measurable functions. The innovations $\{\varepsilon_{t}\}_{t\geq 1}$ are independent and identically distributed. To properly capture important features of the real economy, we allow the supports of $\{Z_{t}\}$ and $\{\varepsilon_{t}\}$ to be vector-valued, and we allow the support of $\{\varepsilon_{t}\}$ to be continuous.
In this setting, $\{Z_{t}\}$ captures both preference shocks and broader idiosyncratic and aggregator factors that jointly shape agents’ intertemporal tradeoffs and consumption and saving behavior. To simplify notation, we denote the partial derivative of utility with respect to consumption by
$u^{\prime}(c,z):=\partial u(c,z)\big/\partial c.$
In addition, we denote $\hat{X}$ as the next period value of a random variable $X$, and
(4) $\mathbbm{E}_{w,z}:=\mathbbm{E}\left[\,\cdot\,\big|\,(w_{0},Z_{0})=(w,z)\right]\quad\text{and}\quad\mathbbm{E}_{z}:=\mathbbm{E}\left[\,\cdot\,\big|\,Z_{0}=z\right].$
2.2. Optimality: Definitions and Fundamental Properties
[…]
Assumption 2.1.
For all $z\in\mathsf{Z}$, $u(\cdot,z)$ is twice differentiable on $(0,\infty)$ with $u^{\prime}(\cdot,z)>0$, $u^{\prime\prime}(\cdot,z)<0$, and $u^{\prime}(c,z)\to\infty$ as $c\to 0$.
Since we did not assume that the utility function is bounded or the discount factor $\beta_{t}$ is less than one, the lifetime utility (1) may not be well defined in $\mathbbm{R}$. To overcome this issue, we define optimality by the overtaking criterion of Brock (1970). Given a feasible policy $c$, we consider the expected sum of utilities over a finite horizon up to time $T$, denoted by
$V_{c,T}(w,z):=\mathbbm{E}_{w,z}\sum_{t=0}^{T}\left(\prod_{i=0}^{t}\beta_{i}\right)u\left(c(w_{t},Z_{t}),Z_{t}\right).$
For two feasible policies $c_{1},c_{2}$, we say that $c_{1}$ overtakes $c_{2}$ if
$\limsup_{T\to\infty}\left[V_{c_{2},T}(w,z)-V_{c_{1},T}(w,z)\right]\leq 0$
[…]
$\hat{w}=\hat{R}(w-c(w,z))+\hat{Y}$.
Noting that $u^{\prime}(\cdot,z)$ is decreasing, the first order condition can be compactly stated as
(5) $u^{\prime}\left(c(w,z),z\right)=\max\left\{\mathbbm{E}_{z}\hat{\beta}\hat{R}u^{\prime}\big(c(\hat{w},\hat{Z}),\hat{Z}\big),u^{\prime}(w,z)\right\}$
[…]
(6) $\lim_{t\to\infty}\mathbbm{E}_{w,z}\left(\prod_{i=0}^{t}\beta_{i}\right)u^{\prime}\left(c(w_{t},Z_{t}),Z_{t}\right)w_{t}=0.$
The following result demonstrates that the first order and transiversality conditions are sufficient for optimality. The proof is similar to Proposition 15.2 and Theorem 15.3 of Toda (2024) and thus omitted.
Proposition 2.1 (Sufficiency of first order and transversality conditions).
If Assumption 2.1 holds, then every feasible policy satisfying the first order and transversality conditions is an optimal policy.
2.3. Existence and Computability of Optimal Consumption
We now study the existence, uniqueness, and computability of a feasible policy satisfying the first order condition (5). We set aside the discussion of the transversality condition (6), as it requires additional assumptions and the proof follows essentially the same steps as Proposition 2.2 in Ma et al. (2020).
[…]
We pair $\mathscr{C}$ with a distance as follows: For each $c_{1},c_{2}\in\mathscr{C}$, define
[…]
which evaluates the maximal difference in terms of marginal utility. While elements of $\mathscr{C}$ are not generally bounded, $\rho$ is a valid metric on $\mathscr{C}$. In particular, $\rho$ is finite on $\mathscr{C}$ since the triangular inequality implies that $\rho(c_{1},c_{2})\leq\left\|u^{\prime}\circ c_{1}-u^{\prime}\right\|+\left\|u^{\prime}\circ c_{2}-u^{\prime}\right\|$, where $\|\cdot\|$ is the standard supremum norm, and the last two terms are finite by (7). In Appendix B, we show that $(\mathscr{C},\rho)$ is a complete metric space.
[…]
(8) $u^{\prime}\left(\xi,z\right)=\max\left\{\mathbbm{E}_{z}\hat{\beta}\hat{R}u^{\prime}\left(c\big(\hat{R}(w-\xi)+\hat{Y},\hat{Z}\big),\hat{Z}\right),u^{\prime}\big(w,z\big)\right\}.$
To show that $T$ is a well defined self-map on $\mathscr{C}$ and characterize the optimal policy via this operator, we impose several key assumptions. Let P(z,z^)P($\hat{z}$) denote the one-step transition probability from $z$ to $\hat{z}$. For $\theta\in\mathbbm{R}$, we define the matrix $K(\theta)$ as follows. For each $z,\hat{z}\in\mathsf{Z}$, let
[…]
where $\pi$ is the probability distribution of $\{\varepsilon_{t}\}$. [^7] The matrix $K(\theta)$ is expressed as a function on $\mathsf{Z}\times\mathsf{Z}$ in (9) but can be represented in traditional matrix notation by enumerating $\mathsf{Z}$. For a square matrix $A$, we use $r(A)$ to denote its spectral radius, defined as
$r(A):=\max\{|\lambda|:\lambda\text{ is an eigenvalue of }A\}.$
In other words, $r(A)$ is the largest absolute value of all its eigenvalues.
Assumption 2.2.
The following conditions hold:
  1. (1)
[…]
  2. (2)
    $r(K(1))<1$, where the matrix $K(\theta)$ is defined by (9).
Following Ma et al. (2020), $\ln r(K(1))$ can be interpreted as the the asymptotic growth rate of average discounted gross payoff on assets, expressed in present-value terms. Accordingly, Assumption 2.2 (2) requires this rate to be negative to ensure that wealth does not diverge asymptotically. This does not preclude the possibility that $\beta_{t}R_{t}\geq 1$ at any given $t$. Rather, it substantially generalizes the standard assumption $\beta R<1$ in the classical optimal savings problem, where both $R_{t}\equiv R$ and $\beta_{t}\equiv\beta$ are constant.
The following theorem establishes the existence and uniqueness of a candidate policy that satisfies the first order optimality condition.
Theorem 2.1 (Existence, uniqueness, and computability of optimal policies).
If Assumptions 2.12.2 hold, then the following statements are true:
  1. (1)
    $T:\mathscr{C}\to\mathscr{C}$ is well defined and has a unique fixed point $c^{*}\in\mathscr{C}$.
  2. (2)
    For all $c\in\mathscr{C}$, we have $\rho(T^{k}c,c^{*})\to 0$ as $k\to\infty$.
Part (2) establishes that, under our assumptions, the familiar time iteration algorithm is globally convergent, provided the initial policy belongs to the candidate class $\mathscr{C}$. Since we set aside the discussion of the transversality condition, which is orthogonal to the main focus of this paper and can be readily incorporated when necessary, we will henceforth refer to $c^{*}$ in Theorem 2.1 as the optimal consumption function.
2.4. Fundamental Properties of Consumption and Saving
We next examine the properties of the optimal consumption function characterized in Theorem 2.1. Throughout, Assumptions 2.12.2 are maintained. The following two propositions establish the monotonicity of the consumption function.
Proposition 2.2 (Monotonicity with respect to wealth).
The optimal consumption and savings functions $c^{*}(w,z)$ and $i^{*}(w,z):=w-c^{*}(w,z)$ are increasing in $w$.
Proposition 2.3 (Monotonicity with respect to income).
If $\{Y_{1t}\}$ and $\{Y_{2t}\}$ are two income processes satisfying $Y_{1t}\leq Y_{2t}$ for all $t$ and $c_{1}^{*}$ and $c_{2}^{*}$ are the corresponding optimal consumption functions, then $c_{1}^{*}\leq c_{2}^{*}$ pointwise on $S\SS$.
The next proposition demonstrates that the borrowing constraint is binding if and only if wealth is below a certain threshold.
Proposition 2.4 (Threshold for saving decision).
For all $c\in\mathscr{C}$, there exists a threshold $\bar{w}_{c}(z)$ such that $Tc(w,z)=w$ if and only if $w\leq\bar{w}_{c}(z)$. In particular, letting
[…]
we have $c^{*}(w,z)=w$ if and only if $w\leq\bar{w}(z)$.
3. Asymptotic Properties of Consumption and Saving
In this section, we study the asymptotic properties of consumption and saving when there are risk aversion shocks. Throughout we assume that Assumptions 2.12.2 hold.
3.1. Setups
We depart from the standard CRRA specification by allowing the level of relative risk aversion to vary with the state, and specify the utility function as
(10) $u(c,z)=\begin{cases}\frac{c^{1-\gamma(z)}}{1-\gamma(z)},&\text{if }\gamma(z)>0\text{ and }\gamma(z)\neq 1,\\ \log c,&\text{if }\gamma(z)=1,\end{cases}$
where $\gamma(z)$ is a state-dependent measure of relative risk aversion. This formulation captures the idea that individuals become more or less tolerant of risk depending on idiosyncratic or aggregate shocks.
Assume that the Markov process $\{Z_{t}\}$ can be decomposed as $Z_{t}=(\bar{Z}_{t},\tilde{Z}_{t})$, where $\{\bar{Z}_{t}\}$ and $\{\tilde{Z}_{t}\}$ are independent Markov processes taking values in $\bar{\mathsf{Z}}:=\{\bar{z}_{1},\dots,\bar{z}_{N}\}$ and $\tilde{\mathsf{Z}}:=\{\tilde{z}_{1},\dots,\tilde{z}_{M}\}$, with transition matrices $\bar{P}=(\bar{p}_{ii^{\prime}})_{1\leq i,i^{\prime}\leq N}$ and $\tilde{P}=(\tilde{p}_{jj^{\prime}})_{1\leq j,j^{\prime}\leq M}$, respectively. In this setting, the transition probability of $\{Z_{t}\}$ satisfies $P=\bar{P}\otimes\tilde{P}$, where $\otimes$ denotes the Kronecker product. In addition, assume that the level of relative risk aversion is driven solely by $\{\bar{Z}_{t}\}$ and satisfies
$0<\gamma(\bar{z}_{1})<\dots<\gamma(\bar{z}_{N}).$
With slight abuse of notation, we denote $\gamma_{i}:=\gamma(\bar{z}_{i})$ for $i=1,\dots,N$. For fixed $i$, we define the $M\times M$ matrix $Q_{i}$ with entries
$Q_{i}(j,k)=\mathbbm{E}_{z_{ij},z_{ik}}\beta(z_{ij},z_{ik},\hat{\varepsilon})R(z_{ij},z_{ik},\hat{\varepsilon})^{1-\gamma_{i}},$
where $z_{ij}=(\bar{z}_{i},\tilde{z}_{j})$, and the expectation is taken with respect to $\hat{\varepsilon}$, conditional on $(Z,\hat{Z})=(z_{ij},z_{ik})$. That is, the elements of $Q_{i}$ represent the corresponding expected future values, conditional on current and future realizations of the exogenous state. We then set
(11) $G_{i}:=\bar{p}_{ii}(\tilde{P}\circ Q_{i}),$
where $\tilde{P}\circ Q_{i}$ represents the Hadamard product (i.e., entry-wise multiplication) of $\tilde{P}$ and $Q_{i}$. Hence $G_{i}$ is obtained by multiplying each entry of $\tilde{P}\circ Q_{i}$ by $\bar{p}_{ii}$.
Recall that a square matrix $A$ is reducible if there is a permutation matrix $B$ such that $B^{\prime}AB$ is block upper triangular with at least two blocks:
$B^{\prime}AB=\begin{bmatrix}A_{11}&A_{12}\\ 0&A_{22}\end{bmatrix}$
where $A_{11}$ and $A_{22}$ are square matrices of size at least one. A square matrix $A$ is called irreducible if it is not reducible. Moreover, recall that $r(A)$ denotes the spectral radius of $A$.
3.2. The Asymptotic MPCs
[…]
Theorem 3.1 (Zero asymptotic MPCs).
Fix $i\in\{1,\dots,N\}$. Suppose either
  1. (1)
    $\sum_{\ell=1}^{i-1}\bar{p}_{i\ell}>0$ and $\mathbbm{P}_{(z,\hat{z})}(\hat{\beta}\hat{R}>0)>0$ for all $z,\hat{z}\in\mathsf{Z}$, or
  2. (2)
    $\bar{p}_{ii}>0$, $G_{i}$ is irreducible, and $r(G_{i})\geq 1$.
Then we have
(12) $\lim_{w\to\infty}\frac{c^{*}(w,z_{ij})}{w}=0\quad\text{for all }\,j=1,\dots,M.$
Theorem 3.1 implies that, if there is a positive probability of transitioning to a lower risk-aversion state and $\beta R>0$ with positive probability at each exogenous state, then the asymptotic MPC is always zero. In applications, the transition matrix often has positive entries, which leads to the following useful result.
Theorem 3.2 (Zero asymptotic MPCs under positive transition probability).
If every entry of $\bar{P}$ is positive and $\mathbbm{P}_{(z,\hat{z})}(\hat{\beta}\hat{R}>0)>0$ for all $z,\hat{z}\in\mathsf{Z}$, then
(13) $\lim_{w\to\infty}\frac{c^{*}(w,z_{ij})}{w}=0\quad\text{for all }z_{ij}\in\mathsf{Z}\text{ with }i\neq 1.$
If in addition $G_{1}$ is irreducible and $r(G_{1})\geq 1$, then (13) holds for all $z_{ij}\in\mathsf{Z}$.
Theorems 3.13.2 above imply that, under mild assumptions, the asymptotic MPC can be strictly positive only if the probability of transitioning to a lower risk-aversion state is zero. In addition, for the asymptotic MPCs to remain positive across all states, the transition matrix of risk aversion must be upper triangular—an exceptional case that rarely arises in practice, as it would require agents to only become (weakly) more risk averse over time. These observations are formalized in the next result. The proof is straightforward and thus omitted.
Proposition 3.1 (Impossibility of positive asymptotic MPCs without downward transitions).
If $\mathbbm{P}_{(z,\hat{z})}(\hat{\beta}\hat{R}>0)>0$ for all $z,\hat{z}\in\mathsf{Z}$, and there exists $z_{ij}\in\mathsf{Z}$ such that
(14) $\lim_{w\to\infty}\frac{c^{*}(w,z_{ij})}{w}>0,$
then $\bar{p}_{i1}=\cdots=\bar{p}_{i,i-1}=0$. In particular, if for all $i>1$, there exists $j$ and $z_{ij}\in\mathsf{Z}$ such that (14) holds, then $\bar{P}$ is upper triangular.
For completeness, we proceed to characterize the complementary cases, in which downward transitions in risk aversion at a specific state occur with zero probability (i.e., $\sum_{\ell=1}^{i-1}\bar{p}_{i\ell}=0$), although such cases are less common in practice.
For each $i\in\{1,\dots,N\}$ and $x\in[1,\infty)^{M}$, we define
(15) $(F_{i}x)(\tilde{z}_{j}):=\left(1+(G_{i}x)(\tilde{z}_{j})^{1/\gamma_{i}}\right)^{\gamma_{i}},\qquad j\in\{1,\dots,M\}.$
Theorem 3.3 (Positive asymptotic MPCs under strictly persistent risk aversion).
Given $i\in\{1,\dots,N\}$, if $\bar{p}_{ii}=1$ and $r(G_{i})<1$, then $F_{i}$ in (15) has a unique fixed point $x_{i}^{*}=(x_{i}^{*}(\tilde{z}_{j}))_{j=1}^{M}$ in $1,\infty)^{M}$ and
(16) $\lim_{w\to\infty}\frac{c^{*}(w,z_{ij})}{w}=x_{i}^{*}(\tilde{z}_{j})^{-1/\gamma_{i}}>0\quad\text{for all }\,j=1,\dots,M.$
Remark 3.1.
Theorem 3.3 is a generalization of the first claim of Theorem 3 in Ma and Toda (2021), in which the level of risk aversion is constant.
Theorem 3.4 (Positive asymptotic MPCs under upward transitions).
Given $i\in\{1,\dots,N\}$, if $\sum_{\ell=1}^{i}\bar{p}_{i\ell}=0$ and there exists $m>0$ such that $R\geq m$ with probability one, then
$\lim_{w\to\infty}\frac{c^{*}(w,z_{ij})}{w}=1\quad\text{for all }\,j=1,\dots,M.$
Theorem 3.4 has important implications. In the classical optimal savings model in which the utility function belongs to the hyperbolic absolute risk aversion (HARA) class, the consumption function is generally concave in wealth (Carroll and Kimball, 1996), even in the presence of both capital income risk and stochastic discounting (Ma et al., 2020). However, Theorem 3.4 shows that this property does not hold when risk aversion shocks are present. This is demonstrated in the following proposition.
Proposition 3.2 (Nonconcave consumption policy).
Given $i\in\{1,\dots,N\}$, if the assumptions of Theorem 3.4 hold, then $c^{*}(w,z_{ij})$ is not concave in $w$ for all $j\in\{1,\dots,M\}$.
Theorems 3.33.4 address the complementary cases of Theorem 3.1, corresponding to $\sum_{\ell=1}^{i-1}\bar{p}_{i\ell}=0$ with $\bar{p}_{ii}=1$ and $\bar{p}_{ii}=0$, respectively. Our final result covers the remaining case with $\sum_{\ell=1}^{i-1}\bar{p}_{i\ell}=0$ and $\bar{p}_{ii}\in(0,1)$. Together, these results exhaust all possible configurations of the transition probability $\bar{P}$.
Theorem 3.5 (Positive asymptotic MPCs under persistent risk aversion).
Given $i\in\{1,\dots,N\}$, if (1) $\sum_{\ell=1}^{i-1}\bar{p}_{i\ell}=0$, (2) $\bar{p}_{ii}\in(0,1)$ and $r(G_{i})<1$, and (3) there exists $m,b>0$ such that $R\geq m$ and $Y\geq b$ with probability one, then (16) holds.
To better highlight our key results, we provide a summary of them in Table 1.
In what follows, we consider some examples, in which for each $z\in\mathsf{Z}$, we denote
$\bar{c}(z):=\lim_{w\to\infty}\frac{c^{*}(w,z)}{w}.$
Example 3.1 (Lower triangular transition matrix).
Suppose $\beta R>0$ with positive probability at each state and that the transition matrix $\bar{P}$ is lower triangular, implying that risk aversion never increases in the future. Then, by Theorem 3.1, we have $\bar{c}(z_{ij})=0$ for all $i>1$ and all $j$. For $i=1$, if $G_{1}$ is irreducible with $r(G_{1})\geq 1$, then $\bar{c}(z_{1j})=0$ for all $j$, and if instead $r(G_{1})<1$, then $\bar{c}(z_{1j})\in(0,1]$ for all $j$.
Table 1. The Asymptotic MPCs
$\sum_{\ell=1}^{i-1}\bar{p}_{i\ell}=0$
Key Assumptions $\sum_{\ell=1}^{i-1}\bar{p}_{i\ell}>0$ or $\bar{p}_{ii}>0$; $r(G_{i})>1$ $\bar{p}_{ii}=1$ $\bar{p}_{ii}=0$ $0<\bar{p}_{ii}<1$
Interpretation Less risk averse with positive probability Generalized knife-edge condition Same risk aversion with probability 1 Strictly more risk averse with probability 1 Weakly more risk averse with positive probability
Other Assumptions $\mathbbm{P}_{(z,\hat{z})}(\hat{\beta}\hat{R}>0)>0$ for all $z,\hat{z}\in\mathsf{Z}$ $G_{i}$ is irreducible $r(G_{i})<1$ $R_{t}\geq m$ for some $m>0$ $r(G_{i})<1$; $R_{t}\geq m$, $Y_{t}\geq b$ for some $m,b>0$
Asymptotic MPCs 0 0 $(0,1]$ 1 $(0,1]$
References Theorem 3.1 Theorem 3.1 Theorem 3.3 Theorem 3.4 Theorem 3.5
Note: $\bar{p}_{ij}$ equals the probability of transitioning from risk aversion state $\gamma_{i}$ to $\gamma_{j}$ in one period. The matrix $G_{i}$ is defined in (11) and $r(G_{i})$ represents its spectral radius.
Example 3.2 (Persistent downward transitions).
Suppose $\beta R>0$ with positive probability at each state and $\bar{p}_{i,i-1}>0$ for all $i>2$, implying that risk aversion follows a persistent Markov chain with strictly positive probability of downward transitions between adjacent states. Then, by Theorem 3.1, we have $\bar{c}(z_{ij})=0$ for all $i>1$ and all $j$. For $i=1$, if $G_{1}$ is irreducible with $r(G_{1})\geq 1$, then $\bar{c}(z_{1j})=0$ for all $j$, and if instead $r(G_{1})<1$, then $\bar{c}(z_{1j})\in(0,1]$ for all $j$.
Example 3.3 (Strictly upward transitions).
Suppose risk aversion increases with probability one in the next period at some state, and $R_{t}$ has a strictly positive lower bound (which can be arbitrarily close to zero). Then, by Theorem 3.4, the asymptotic MPC at that state equals one, and by Proposition 3.2, the consumption function is not concave in wealth.
Example 3.4 (Constant returns without capital income risk).
Suppose the gross rate of return $R_{t}$ is constant at $R$. By Proposition 8 of Ma and Toda (2021), in the absence of preference shocks, a zero asymptotic MPC can occur only when $R<1$. This is empirically implausible, since it implies a negative rate of return to wealth. In contrast, when preference shocks are present, Theorem 3.1 above demonstrates that the asymptotic MPC is zero (except at the lowest risk aversion state) whenever the probability of transitioning to a lower risk aversion state and the probability that $\beta R>0$ are positive, without any further restriction on $R$. In particular, zero asymptotic MPCs can arise even when $R\geq 1$.
Example 3.5 (Classical optimal savings problem with preference shocks).
Consider the classical optimal savings problem, where $R_{t}\equiv R$ and $\beta_{t}\equiv\beta$ are positive constants with $\beta R<1$ and $Y_{t}$ has a positive lower bound. Without preference shocks, Ma and Toda (2021) show that the asymptotic MPC is
$\bar{c}(z)=\begin{cases}0,&\text{if $\beta R^{1-\gamma}\geq 1$,}\\ 1-(\beta R^{1-\gamma})^{1/\gamma},&\text{otherwise.}\end{cases}$
With preference shocks, Theorems 3.13.5 imply that the asymptotic MPC is
$\bar{c}(z_{ij})=\begin{cases}0,&\text{if $\sum_{\ell=1}^{i-1}\bar{p}_{i\ell}>0$ or $\bar{p}_{ii}\beta R^{1-\gamma_{i}}\geq 1$,}\\ 1-(\bar{p}_{ii}\beta R^{1-\gamma_{i}})^{1/\gamma_{i}},&\text{otherwise.}\end{cases}$
In particular, the asymptotic MPC is zero whenever there is a positive probability of transitioning from $\gamma_{i}$ to a lower risk aversion state. Hence, the set of parameter values leading to zero asymptotic MPCs is substantially broadened: zero asymptotic MPCs can arise even when $\beta R^{1-\gamma}<1$. Moreover, Theorem 3.4 and Proposition 3.2 imply that the asymptotic MPC equals one and the consumption function is not concave in wealth at state $\gamma_{i}$ when $\sum_{\ell=1}^{i}\bar{p}_{i\ell}=0$.
Example 3.6 (Optimal savings with iid return and discounting).
Suppose, as in Benhabib et al. (2015), $\{\beta_{t}\}$ and $\{R_{t}\}$ are iid driven by innovation $\{\varepsilon_{t}\}$ with positive lower bounds, and $\{Y_{t}\}$ has a positive lower bound. Without preference shocks, Ma and Toda (2021) show that
$\bar{c}(z)=\begin{cases}0,&\text{if }\mathbbm{E}\beta R^{1-\gamma}\geq 1\\ 1-(\mathbbm{E}\beta R^{1-\gamma})^{1/\gamma},&\text{otherwise.}\end{cases}$
With preference shocks, Theorems 3.13.5 imply that the asymptotic MPC is
$\bar{c}(z_{ij})=\begin{cases}0,&\text{if $\sum_{\ell=1}^{i-1}\bar{p}_{i\ell}>0$ or $\bar{p}_{ii}\mathbbm{E}\beta R^{1-\gamma_{i}}\geq 1$,}\\ 1-(\bar{p}_{ii}\mathbbm{E}\beta R^{1-\gamma_{i}})^{1/\gamma_{i}},&\text{otherwise.}\end{cases}$
Same as in Example 3.5, the set of parameter values yielding zero asymptotic MPCs is significantly broadened: such outcomes can occur even when $\mathbbm{E}\beta R^{1-\gamma_{i}}<1$. Moreover, the asymptotic MPC equals one and the optimal consumption function is not concave in wealth at state $\gamma_{i}$ when $\sum_{\ell=1}^{i}\bar{p}_{i\ell}=0$.
3.3. The Asymptotic Saving Rate
As mentioned earlier, we follow the literature and define an agent’s saving rate as the change in net worth divided by total income, excluding capital losses:
$s_{t+1}=\frac{w_{t+1}-w_{t}}{\max\{(R_{t+1}-1)(w_{t}-c_{t}),0\}+Y_{t+1}}.$
Based on (2) and some simple algebra, we can show that [^9] Here, we define $x^{+}=\max\{x,0\}$ and $x^{-}=-\min\{x,0\}$ for each $x\in\mathbbm{R}$.
$s_{t+1}=1-\frac{(R_{t+1}-1)^{-}(1-c_{t}/w_{t})+c_{t}/w_{t}}{(R_{t+1}-1)^{+}(1-c_{t}/w_{t})+Y_{t+1}/w_{t}}\in(-\infty,1).$
Letting $w_{t}\to\infty$, the asymptotic saving rate, which can be interpreted as the saving rate of an infinitely wealthy agent, is given by
(17) $\bar{s}:=1-\frac{(\hat{R}-1)^{-}(1-\bar{c})+\bar{c}}{(\hat{R}-1)^{+}(1-\bar{c})}\in[-\infty,1],$
where, as in Section 3.2, $\bar{c}$ denotes the asymptotic MPC (suppose it is well defined).
When the rate of return on wealth is positive (i.e., $R_{t}>1$) and the asymptotic MPC is zero, the asymptotic saving rate converges to 100%. Conversely, if the asymptotic MPC is positive and sufficiently large, the asymptotic saving rate becomes negative, implying that precautionary saving motive vanishes as wealth grows without bound.
4. Quantitative Analysis
In this section, we complement the theory with a quantitative analysis that serves two purposes. First, using a parsimonious quantitative specification, we provide numerical illustrations that render the main theoretical results transparent and economically intuitive. Second, within a more general quantitative environment, we compute generalized impulse response functions to elucidate the economic mechanisms through which preference shocks generate qualitative changes in model dynamics.
4.1. Model Specification
Consider an optimal savings problem with risk aversion shocks. The discount factor and the gross rate of return on wealth are held constant over time with $\beta_{t}\equiv\beta$ and $R_{t}\equiv R$. Risk aversion follows a discretized AR(1) process:
$\gamma_{t+1}=\mu_{\gamma}(1-\rho_{\gamma})+\rho_{\gamma}\gamma_{t}+\sigma_{\gamma}\sqrt{1-\rho_{\gamma}^{2}}\varepsilon_{t+1}^{\gamma},\quad(\varepsilon_{t}^{\gamma})\mathrel{{\mathop{\sim}\limits^{\textrm{ {\sc iid }}}}}N(0,1).$
This specification is a special case of our theory, with the preference state given by $\bar{Z}_{t}=\gamma_{t}$. For nonfinancial income, we allow for alternative specifications depending on the quantitative experiment under consideration. The model is solved using the endogenous grid method of Carroll (2006). Detailed implementation and algorithmic steps are provided in Appendix E.
In our quantitative experiments, we work at an annual frequency and set $\beta=0.95$ and $R=1.02$. In our baseline setting, risk aversion is parameterized by $\rho_{\gamma}=0.6$, $\mu_{\gamma}=4$, and $\sigma_{\gamma}=1$, and discretized into a finite-state Markov chain using the method of Tauchen (1986). The number of discrete states is allowed to vary across experiments. In particular, for the impulse response analysis in Section 4.3, we employ a fine discretization to improve approximation accuracy. In all specifications, the minimum, median (and mean), and maximum values of $\gamma_{t}$ are $1$, 44, and $7$, respectively, which lie well within the range commonly used in empirical research.
4.2. The Asymptotic Saving Rates
To better illustrate the implications of our theory, we begin by comparing our model with the standard optimal savings problem without preference shocks. To this end, we consider simply $Y_{t}\equiv 1$ and discretize $\{\gamma_{t}\}$ into a three-state Markov chain. The resulting state space and transition matrix are
$\gamma_{t}\in\{1,4,7\}\quad\text{and}\quad\bar{P}=\begin{bmatrix}0.6416&0.3538&0.0001\\ 0.0303&0.9392&0.0305\\ 0.0001&0.3538&0.6461\end{bmatrix}$.
The stationary mean of this discrete process coincides with the median state $\gamma_{t}=4$. At this state, the probability of remaining at $\gamma_{t}=4$ next period is 93.92%, while the probability of transitioning to the lower risk aversion state $\gamma_{t}=1$ is only 3.03%.
To ensure a fair comparison, we contrast the optimal consumption function when $\gamma_{t}$ is fixed at its stationary mean $\mathbbm{E}\gamma_{t}=4$ in our stochastic risk aversion model with that of the standard optimal savings problem featuring CRRA utility with a constant risk aversion coefficient $\gamma=4$. Figure 1 plots the results. [^10] The main difference between the left and right panels lies in the range of wealth displayed. Moreover, because the right panel is set on a log scale, the consumption function appears nonconcave, though it is in fact concave. A log scale is similarly used in Figure 2 when required. Despite the fact that current risk aversion is identical across the two models and that there is a high probability of remaining at the same level of risk aversion next period, consumption is substantially lower under stochastic risk aversion (red dotted line) than under constant risk aversion (blue solid line), reflecting the precautionary behavior induced by the possibility of future changes in risk aversion.
Figure 2 plots the consumption rate, defined as $c^{*}(w,z)/w$, and the saving rate as a function of wealth. The left panel shows that, in the standard optimal savings model, the consumption rate (blue solid line) converges to the theoretical asymptotic MPC, which is strictly positive and equal to $1-(\beta R^{1-\gamma})^{1/\gamma}=0.0273$ as derived in Example 3.5. By contrast, in the stochastic risk aversion model, the consumption rate exhibits a pronounced downward trend and continues to decline even at very high levels of wealth. This pattern is fully consistent with Theorem 3.1, which establishes that the asymptotic MPC is equal to zero.
Figure 1. The optimal consumption function
Figure 2. The consumption and saving rates
The right panel plots the corresponding saving rates. In the standard optimal savings model, saving rates are negative and converge to their theoretical asymptotic value of $40.31\%$, as given by (17), in stark contradiction with the empirical evidence on wealthy households. In contrast, under stochastic risk aversion, the saving rate converges to its theoretical limit of $100\%$. Taken together, these results illustrate that even minor probabilities of shifts in risk aversion can have sizable effects on the saving behavior, suggesting that stochastic risk aversion offers a promising mechanism for generating persistently high saving rates among wealthy households, in line with the patterns observed in the data.
4.3. Impulse Response Analysis
In this section, we systematically examine how changes in risk aversion influence consumption and saving behavior using impulse response functions (IRFs). To capture the model’s nonlinear dynamics, we construct IRFs as state-dependent random variables, following Koop et al. (1996). These IRFs allow us to trace how shocks to risk aversion propagate over time and affect consumption, saving rates, and consumption volatility across different states. [^11] Consumption volatility in period-$t$ is defined as the conditional variance: $\operatorname{Var}_{t-1}[c^{*}(w_{t},Z_{t})]:=\mathbbm{E}_{t-1}[c^{*}(w_{t},Z_{t})]^{2}-[\mathbbm{E}_{t-1}c^{*}(w_{t},Z_{t})]^{2}.$
We follow the standard earnings specification in the literature and assume that labor income evolves according to
$\log Y_{t}=\eta_{t}+\nu_{t},\quad(\nu_{t})\sim N(0,\sigma_{v}^{2}),$
$\eta_{t+1}=\rho_{\eta}\eta_{t}+\varepsilon_{t}^{\eta},\quad(\varepsilon_{t}^{\eta})\sim N(0,\sigma_{\eta}^{2}).$
Consistent with the empirical estimates reported in Lander (2024), we set $\rho_{\eta}=0.9735$, $\sigma_{\eta}=0.1272$, and $\sigma_{\nu}=0.2102$, based on PSID data over 1970-2012.
To ensure sufficient numerical accuracy, we discretize $\{\gamma_{t}\}$ into a 121-state Markov chain following Tauchen (1986). In addition, the persistent income component $\{\eta_{t}\}$ is discretized into 5 states, while expectations over the transitory shock $\{\nu_{t}\}$ are computed using a 7-point Gauss–Hermite quadrature. We also conducted experiments using the discretization method of Rouwenhorst (1995). Sensitivity analysis confirms that the results are robust to both parametric choices and discretization methods.
Figure 3. Impulse responses to a positive one-unit risk aversion shock
Figure 3 reports impulse response functions to a one-unit increase in risk aversion at the beginning of period $t$. As noted earlier, the IRFs are computed as state-dependent responses, conditional on different initial levels of wealth, risk aversion, and the persistent income component $(w_{t-1},\gamma_{t-1},\eta_{t-1})$. All experiments are conducted conditional on $\eta_{t-1}=0$, its stationary mean.
The top row of Figure 3 illustrates heterogeneity with respect to initial wealth, holding risk aversion fixed at its stationary mean $\gamma_{t-1}=4$. The top-left panel shows that an increase in risk aversion induces an immediate and persistent decline in consumption, with the magnitude of the response rising sharply in wealth. When wealth rises from $w_{t-1}=5$ to $10$ and $50$, contemporaneous consumption falls by 4.67%, 7.89%, and 20.23%, respectively, with effects lasting roughly 2–6 years. This nonlinearity reflects the fact that low-wealth households are constrained in their ability to adjust savings, so consumption responds weakly to preference shocks, whereas wealthy households can reallocate intertemporal resources more flexibly.
The top-middle panel shows that this adjustment operates primarily through saving. A one-unit increase in risk aversion raises the saving rate by 5.81%, 10.15%, and 28.05% at $w_{t-1}=5,10,$ and $50$, respectively, with effects persisting for 4–8 years. Thus, as wealth rises, the response shifts from limited consumption adjustment toward substantial increases in saving. This pattern aligns with our theoretical mechanism: at high wealth levels, precautionary motives dominate, prompting agents to accumulate assets aggressively even in response to transitory risk aversion shocks. The top-right panel further shows that consumption volatility declines following a positive risk aversion shock, particularly for wealthier households, reflecting stronger consumption smoothing driven by heightened precautionary behavior.
The middle row of Figure 3 illustrates heterogeneity with respect to initial risk aversion, holding initial wealth fixed at $w_{t-1}=10$. When initial risk aversion is high ($\gamma_{t-1}=6$), the same shock generates relatively modest responses: consumption falls by 4.84% and the saving rate rises by 5.53%. By contrast, when agents start from a low risk aversion state ($\gamma_{t-1}=1$), consumption drops sharply by 19.26% and saving increases by 32.21%. This asymmetry reflects state-dependent amplification: a positive shock represents a larger proportional change in effective prudence when baseline risk aversion is lower, triggering a stronger reoptimization of intertemporal choices. Consistent with this mechanism, the decline in consumption volatility is also largest at low initial risk aversion.
The bottom row of Figure 3 examines the role of the persistence of risk aversion shocks, governed by $\rho_{\gamma}$, holding initial wealth and risk aversion fixed at $(w_{t-1},\gamma_{t-1})=(10,4)$. Higher persistence markedly amplifies and prolongs the responses of consumption and saving. As $\rho_{\gamma}$ increases, households cut consumption more sharply, raise saving more persistently, and experience larger declines in consumption volatility. Intuitively, greater persistence raises the expected duration of elevated risk aversion, strengthening precautionary motives and leading agents to treat preference shocks as quasi-permanent. This persistence channel thus reinforces the state-dependent and nonlinear dynamics emphasized by our theory.
Taken together, these impulse responses deliver three key insights. First, shocks to risk aversion generate economically large and persistent movements in consumption and saving. Second, these effects are highly state dependent, varying systematically with households’ wealth, the current level of risk aversion, and the persistence of preference shocks. Third, the IRFs provide a quantitative illustration of the core theoretical mechanism: stochastic risk preferences alone can fundamentally reshape saving behavior, even in environments with constant returns and discounting, accounting for the high and persistent saving rates observed among wealthy households.
Appendix A Preliminaries
Let $(\Omega,\mathscr{F},\mathbbm{P})$ be a fixed probability space on which all random variables are defined, and let $\mathbbm{E}$ denote expectations under $\mathbbm{P}$. The state process $\{Z_{t}\}$ and the innovation process $\{\varepsilon_{t}\}$ introduced in (3) are defined on this space, with $\pi$ denoting the marginal distribution of $\{\varepsilon_{t}\}$. Let $\{\mathscr{F}_{t}\}$ be the natural filtration generated by $\{Z_{t}\}$ and $\{\varepsilon_{t}\}$. For $z\in\mathsf{Z}$, we write $\mathbbm{P}_{z}$ for probability conditional on $Z_{0}=z$ and $\mathbbm{E}_{z}$ for expectation under $\mathbbm{P}_{z}$.
For a stochastic process $\varphi_{t}=\varphi(Z_{t-1},Z_{t},\varepsilon_{t})$, where $\varphi$ is a nonnegative measurable function, we define the matrix $L_{\varphi}$ as
[…]
The matrix $L_{\varphi}$ is expressed as a function on $\mathsf{Z}\times\mathsf{Z}$ in (18), but it can be represented in traditional matrix notation by enumerating $\mathsf{Z}$. Recall that $\mathbbm{R}^{\mathsf{Z}}$ is the set of real-valued functions on $\mathsf{Z}$. In what follows, we treat $L_{\varphi}$ as a linear operator on $\mathbbm{R}^{\mathsf{Z}}$. A proof by induction shows that, for any $h\in\mathbbm{R}^{\mathsf{Z}}$, we have
(19) $(L_{\varphi}^{n}\,h)(z)=\mathbbm{E}_{z}\prod_{t=1}^{n}\varphi_{t}h(Z_{t}),$
where $L_{\varphi}^{n}$ is the $n$-th composition of the operator $L_{\varphi}$ with itself, or equivalently, the $n$-th power of the matrix $L_{\varphi}$. The following result holds. The proof is similar to the proof of Lemma 4.1 of Ma et al. (2020) and thus omitted.
Lemma A.1.
Let $\{\varphi_{t}\}$ and $L_{\varphi}$ be as defined in (18). If $r(L_{\varphi})<1$, then there exist $N\in\mathbbm{N}$ and $\sigma\in(r(L_{\varphi}),1)$ such that $\max_{z\in\mathsf{Z}}\mathbbm{E}_{z}\prod_{t=1}^{n}\varphi_{t}<\sigma^{n}$ for all $n\geq N$.
Appendix B Proof of Section 2 Results
Throughout this section, we assume that Assumptions 2.12.2 hold. Without loss of generality, we may assume $\mathsf{Z}=\{1,\dots,Z\}$ in the related proofs of this section.
Proposition B.1.
$(\mathscr{C},\rho)$ is a complete metric space.
Proof.
The proof is a straightforward extension of Proposition 4.1 of Li and Stachurski (2014) and thus omitted. A full proof is available from the authors upon request. ∎
Proposition B.2.
[…]
Proof.
For given $c\in\mathscr{C}$, we can rewrite (8) as
(20) $u^{\prime}(\xi,z)=\psi_{c}(\xi,w,z):=\max\{g_{c}(\xi,w,z),u^{\prime}(w,z)\},$
where $g_{c}$ is a function on
[…]
defined by
[…]
To establish existence, we apply the intermediate value theorem. It suffices to verify the following three conditions:
  1. (a)
    the map $\xi\mapsto\psi_{c}(\xi,w,z)$ is continuous on $(0,w]$,
  2. (b)
    there exists a $\xi\in(0,w]$ such that $u^{\prime}(\xi,z)\geq\psi_{c}(\xi,w,z)$, and
  3. (c)
    there exists a $\xi\in(0,w]$ such that $u^{\prime}(\xi,z)\leq\psi_{c}(\xi,w,z)$.
For part (a), it is sufficient to show that $\xi\mapsto g_{c}(\xi,w,z)$ is continuous on $(0,w]$. To this end, fix $\xi\in(0,w]$ and let $\xi_{n}\to\xi$. Since $c\in\mathscr{C}$, there exists a constant $M\in\mathbbm{R}_{+}$ such that
[…]
By the monotonicity of $u^{\prime}$, we have
(23) $\hat{\beta}\hat{R}u^{\prime}[c(\hat{R}\left(w-\xi\right)+\hat{Y},\hat{Z}),\hat{Z}]\leq\hat{\beta}\hat{R}u^{\prime}[c(\hat{Y},\hat{Z}),\hat{Z}]\leq\hat{\beta}\hat{R}u^{\prime}(\hat{Y},\hat{Z})+\hat{\beta}\hat{R}M.$
Furthermore, by Assumption 2.2, we know that, for all $z\in\mathsf{Z}$,
$\mathbbm{E}_{z}\hat{\beta}\hat{R}u^{\prime}(\hat{Y},\hat{Z})<\infty\quad\text{and}\quad\mathbbm{E}_{z}\hat{\beta}\hat{R}$$\infty$.
Hence, by the dominated convergence theorem and the continuity of $c$, we can conclude that $g(\xi_{n},w,z)\to g(\xi,w,z)$. This proves that $\xi\mapsto\psi_{c}(\xi,w,z)$ is continuous.
Part (b) clearly holds, since $u^{\prime}(\xi,z)\to\infty$ as $\xi\to 0$ and $\xi\mapsto\psi_{c}(\xi,w,z)$ is increasing and always finite (since it is continuous as shown in the previous paragraph). Part (c) is also trivial (just set $\xi=w$). ∎
Proposition B.3.
We have $Tc\in\mathscr{C}$ for all $c\in\mathscr{C}$.
Proof.
Fix $c\in\mathscr{C}$ and let $\psi_{c}$ and $g_{c}$ be defined as in (20) and (22), respectively.
Step 1. We show that $Tc$ is continuous. To apply a standard fixed point parametric continuity result, such as Theorem B.1.4 of Stachurski (2009), we first demonstrate that $\psi_{c}$ is jointly continuous on the set $G$ defined in (21). This will hold if $g_{c}$ is jointly continuous on $G$. For any $\{(\xi_{n},w_{n},z_{n})\}$ and $(\xi,w,z)$ in $G$ with $(\xi_{n},w_{n},z_{n})\to(\xi,w,z)$, we need to show that $g_{c}(\xi_{n},w_{n},z_{n})\to g_{c}(\xi,w,z)$. To this end, we define
[…]
By Fatou’s lemma and Theorem 1.1 of Feinberg et al. (2014), we have
[…]
which implies
[…]
This shows that $g_{c}$ is then continuous, since the inequality above is equivalent to the statement
$\liminf_{n\to\infty}g_{c}(\xi_{n},w_{n},z_{n})\geq g_{c}(\xi,w,z)\geq\limsup_{n\to\infty}g_{c}(\xi_{n},w_{n},z_{n}).$
Hence, $\psi_{c}$ is continuous on $G$, as required. Moreover, since $\xi\mapsto\psi_{c}(\xi,w,z)$ takes values in the closed interval
$I(w,z):=\left[u^{\prime}(w,z),u^{\prime}(w,z)+\mathbbm{E}_{z}\hat{\beta}\hat{R}(u^{\prime}(\hat{Y},\hat{Z})+M)\right]$
and the correspondence $(w,z)\mapsto I(w,z)$ is nonempty, compact-valued and continuous, Theorem B.1.4 of Stachurski (2009) implies that $Tc$ is continuous on $S\SS$.
Step 2. We show that $Tc$ is increasing in $w$. Suppose that for some $z\in\mathsf{Z}$ and $w_{1},w_{2}\in(0,\infty)$ with $w_{1}<w_{2}$, we have $\xi_{1}:=Tc(w_{1},z)>Tc(w_{2},z)=:\xi_{2}$. Since $c$ is increasing in $w$ by assumption, $\psi_{c}$ is increasing in $\xi$ and decreasing in $w$. Thus, we have $u^{\prime}(\xi_{1},z)<u^{\prime}(\xi_{2},z)=\psi_{c}(\xi_{2},w_{2},z)\leq\psi_{c}(\xi_{1},w_{1},z)=u^{\prime}(\xi_{1},z)$, which is a contradiction.
[…]
Step 4. Since $u^{\prime}[Tc(w,z),z]\geq u^{\prime}(w,z)$, we have
$\left|u^{\prime}[Tc(w,z),z]-u^{\prime}(w,z)\right|=u^{\prime}[Tc(w,z),z]-u^{\prime}(w,z)$
$\leq\mathbbm{E}_{z}\hat{\beta}\hat{R}u^{\prime}\,[c(\hat{R}(w-Tc(w,z))+\hat{Y},\,\hat{Z}),\hat{Z}\,]\leq\mathbbm{E}_{z}\hat{\beta}\hat{R}(u^{\prime}(\hat{Y},\hat{Z})+M)$
[…]
For each $h:(0,\infty)\to\mathbbm{R}_{+}^{\mathsf{Z}}$, we write
$h(w)=(h_{1}(w),\dots,h_{Z}(w)).$
To prove Theorem 2.1, let $\mathscr{H}$ denote the set of all continuous functions $h:(0,\infty)\to\mathbbm{R}_{+}^{\mathsf{Z}}$ such that each $h_{z}$ is decreasing and $w\mapsto h_{z}(w)-u^{\prime}(w,z)$ is bounded and nonnegative. For any $h\in\mathscr{H}$, we define $(\tilde{T}h)_{z}(w)$ as the value $\kappa$ that solves
(24) $\kappa=\max\left\{\mathbbm{E}_{z}\,\hat{\beta}\hat{R}\,h(\hat{R}\,[w-(u^{\prime}(\cdot,z))^{-1}(\kappa)]+\hat{Y},\hat{Z}),u^{\prime}(w,z)\right\},$
where, with a slight abuse of notation, we denote $h_{z}(w):=h(w,z)$. Moreover, we consider the bijection $H:\mathscr{C}\to\mathscr{H}$ defined by $(Hc)_{z}(w):=u^{\prime}(c(w,z),z)$. The next lemma implies that $\tilde{T}$ is a well-defined self-map on $\mathscr{H}$, as well as topologically conjugate to $T$.
Lemma B.1.
The operator $\tilde{T}\colon\mathscr{H}\to\mathscr{H}$ and satisfies $\tilde{T}H=HT$ on $\mathscr{C}$.
Proof.
[…]
(25) $u^{\prime}(\xi,z)=\max\left\{\mathbbm{E}_{z}\hat{\beta}\hat{R}u^{\prime}(c(\hat{R}\left(w-\xi\right)+\hat{Y},\hat{Z}),\hat{Z}),u^{\prime}(w,z)\right\}.$
We need to show that $HTc$ and $\tilde{T}Hc$ evaluate to the same number at $(w,z)$. In other words, we need to verify that $u^{\prime}(\xi,z)$ is the solution to
$\kappa=\max\left\{\mathbbm{E}_{z}\hat{\beta}\hat{R}u^{\prime}(c(\hat{R}\,[w-u^{\prime}(\cdot,z)^{-1}(\kappa)]+\hat{Y},\hat{Z}),\hat{Z}),u^{\prime}(w,z)\right\}.$
But this follows immediately from (25). Hence, we have shown that $\tilde{T}H=HT$ on $\mathscr{C}$. Since $H\colon\mathscr{C}\to\mathscr{H}$ is a bijection, we have $\tilde{T}=HTH^{-1}$. Moreover, Proposition B.3 ensures that $T\colon\mathscr{C}\to\mathscr{C}$, and hence $\tilde{T}\colon\mathscr{H}\to\mathscr{H}$. This completes the proof. ∎
To prove Theorem 2.1, we apply the Perov contraction theorem following Toda (2021). For $h,g\in\mathscr{H}$ and $z\in\mathsf{Z}$, define
$d_{z}(h,g):=\sup_{w\in(0,\infty)}|h_{z}(w)-g_{z}(w)|.$
Note that $d_{z}$ is always finite by the definition of $\mathscr{H}$. Define the vector-valued metric $d:\mathscr{H}\times\mathscr{H}\to\mathbbm{R}_{+}^{\mathsf{Z}}$ by $d(h,g)=(d_{1}(h,g),\dots,d_{Z}(h,g))$. By Proposition B.1 and the discussion in Section 2 of Toda (2021), $(\mathscr{H},d)$ is a complete vector-valued metric space.
Lemma B.2.
Let $B:=K(1)$. Then $\tilde{T}$$\mathscr{H}\to\mathscr{H}$ is a Perov contraction with coefficient matrix $B$, that is, for any $h_{1},h_{2}\in\mathscr{H}$, we have $d(\tilde{T}h_{1},\tilde{T}h_{2})\leq Bd(h_{1},h_{2})$.
Proof.
By Assumption 2.2 (2), we have $r(B)\in[0,1)$. By Theorem 3 of Toda (2021), it suffices to verify the following two conditions:
  1. (1)
    If $h_{1},h_{2}\in\mathscr{H}$ and $h_{1}\leq h_{2}$ pointwise, then $\tilde{T}h_{1}\leq\tilde{T}h_{2}$ pointwise.
  2. (2)
    For any $h\in\mathscr{H}$ and $a\in\mathbbm{R}_{+}^{\mathsf{Z}}$, we have $\tilde{T}(h+a)\leq\tilde{T}h+Ba$ pointwise.
[…]
$u^{\prime}(\xi_{1},z)$ $=\max\{\mathbbm{E}_{z}\hat{\beta}\hat{R}\,h_{1}(\hat{R}(w-\xi_{1})+\hat{Y},\hat{Z}),u^{\prime}(w,z)\}$
$\leq\max\{\mathbbm{E}_{z}\hat{\beta}\hat{R}\,h_{2}(\hat{R}(w-\xi_{1})+\hat{Y},\hat{Z}),u^{\prime}(w,z)\}$
$\leq\max\{\mathbbm{E}_{z}\hat{\beta}\hat{R}\,h_{2}(\hat{R}(w-\xi_{2})+\hat{Y},\hat{Z}),u^{\prime}(w,z)\}=u^{\prime}(\xi_{2},z),$
where the first inequality uses $h_{1}\leq h_{2}$ and the second inequality uses the fact that $h_{2}$ is decreasing in $w$ and $\xi_{1}<\xi_{2}$. Since $u^{\prime}$ is strictly decreasing, we obtain $\xi_{1}\geq\xi_{2}$, which is a contradiction.
[…]
$(\tilde{T}(h+a))_{z}(w)=\kappa(a)=u^{\prime}(\xi(a),z)$
$=\max\{\mathbbm{E}_{z}\hat{\beta}\hat{R}(h+a)(\hat{R}(w-\xi(a))+\hat{Y},\hat{Z}),u^{\prime}(w,z)\}$
$=\max\{\mathbbm{E}_{z}\hat{\beta}\hat{R}h(\hat{R}(w-\xi(a))+\hat{Y},\hat{Z})+\mathbbm{E}_{z}\hat{\beta}\hat{R}a_{\hat{Z}},u^{\prime}(w,z)\}$
$\leq\max\{\mathbbm{E}_{z}\hat{\beta}\hat{R}h(\hat{R}(w-\xi(a))+\hat{Y},\hat{Z}),u^{\prime}(w,z)\}+\mathbbm{E}_{z}\hat{\beta}\hat{R}a_{\hat{Z}}$
$\leq\max\{\mathbbm{E}_{z}\hat{\beta}\hat{R}h(\hat{R}(w-\xi(0))+\hat{Y},\hat{Z}),u^{\prime}(w,z)\}+\mathbbm{E}_{z}\hat{\beta}\hat{R}a_{\hat{Z}}$
$=(\tilde{T}h)_{z}(w)+(Ba)_{z}.$
Therefore, $\tilde{T}(h+a)\leq\tilde{T}h+Ba$ pointwise. ∎
Proof of Theorem 2.1.
By Lemma B.2 and Theorem 2 of Toda (2021), there exists a unique fixed point $h^{*}\in\mathscr{H}$ of $\tilde{T}$. From the topological conjugacy result in Lemma B.1, we have $\tilde{T}=HTH^{-1}$, so there exists a unique fixed point $c^{*}\in\mathscr{C}$ of $T$, and claim (i) is verified.
To see that claim (2) holds, take any $\sigma\in(r(B),1)$. With a slight abuse of notation, let $\|\cdot\|$ denote the supremum norm on $\mathbbm{R}^{\mathsf{Z}}$, the operator norm it induces on $\mathbbm{R}^{\mathsf{Z}\times\mathsf{Z}}$, as well as the supremum norm on the space of functions from $(0,\infty)$ to $\mathbbm{R}^{\mathsf{Z}}$. Since $\tilde{T}^{n}=HT^{n}H^{-1}$ and $\tilde{T}$ is a Perov contraction with coefficient matrix $B$ with spectral radius $r(B)<1$, it follows from the definition of $\rho$, $\mathscr{C}$, $\mathscr{H}$ and Lemma A.1 that
$\rho(T^{n}c,c^{*})$ $=\rho(T^{n}c,T^{n}c^{*})=\|HT^{n}c-HT^{n}c^{*}\|=\|\tilde{T}^{n}Hc-\tilde{T}^{n}Hc^{*}\|$
$\leq\|B^{n}\|\|Hc-Hc^{*}\|\leq\sigma^{n}\|Hc-Hc^{*}\|$
for large enough $n$. Therefore, $\rho(T^{n}c,c^{*})\to 0$ as $n\to\infty$. ∎
Our next goal is to prove Proposition 2.2. We begin by defining
$\mathscr{C}_{0}=\left\{c\in\mathscr{C}\colon w\mapsto w-c(w,z)\text{ is increasing for all }z\in\mathsf{Z}\right\}.$
Lemma B.3.
$\mathscr{C}_{0}$ is a closed subset of $\mathscr{C}$, and $Tc\in\mathscr{C}_{0}$ for all $c\in\mathscr{C}_{0}$.
Proof.
[…]
Fix $c\in\mathscr{C}_{0}$. We now show that $\xi:=Tc\in\mathscr{C}_{0}$. By Proposition B.3, $\xi\in\mathscr{C}$, hence it suffices to show that $w\mapsto w-\xi(w,z)$ is increasing. Suppose not. Then there exist $z\in\mathsf{Z}$ and $w_{1},w_{2}\in(0,\infty)$ such that $w_{1}<w_{2}$ and $w_{1}-\xi(w_{1},z)>w_{2}-\xi(w_{2},z)$. Since $w_{1}-\xi(w_{1},z)\geq 0$, $w_{2}-\xi(w_{2},z)\geq 0$ and $\xi(w_{1},z)\leq\xi(w_{2},z)$ by Proposition B.3, we must have $\xi(w_{1},z)<w_{1}$ and $\xi(w_{1},z)<\xi(w_{2},z)$. However, the definition of $T$ combined with the concavity of $u$ then implies that
[…]
$\leq\mathbbm{E}_{z}\hat{\beta}\hat{R}u^{\prime}(c(\hat{R}\,[w_{2}-\xi(w_{2},z)]+\hat{Y},\hat{Z}),\hat{Z})\leq u^{\prime}(\xi(w_{2},z),z),$
which gives $\xi(w_{1},z)\geq\xi(w_{2},z)$, yielding a contradiction. Hence, $w\mapsto w-\xi(w,z)$ is increasing and $T$ is a self-map on $\mathscr{C}_{0}$. ∎
Proof of Proposition 2.2.
Since $T$ maps elements of the closed subset $\mathscr{C}_{0}$ into itself by Lemma B.3, Theorem 2.1 implies that $c^{*}\in\mathscr{C}_{0}$. Hence, the stated claims hold. ∎
Proof of Proposition 2.3.
Let $T_{j}$ be the time iteration operator for the income process $j$ as defined in Proposition B.3. It suffices to show $T_{1}c\leq T_{2}c$ for all $c\in\mathscr{C}$. To see this, by the monotonicity of $T_{j}$, we have $T_{j}c_{1}\leq T_{j}c_{2}$ whenever $c_{1}\leq c_{2}$. Thus, if $T_{1}c\leq T_{2}c$ for all $c\in\mathscr{C}$, then for any $c_{1},c_{2}\in\mathscr{C}$ with $c_{1}\leq c_{2}$, $T_{1}c_{1}\leq T_{1}c_{2}\leq T_{2}c_{2}$. Iterating from any $c\in\mathscr{C}$ and using Theorem 2.1, we obtain $c_{1}^{*}=\lim_{n\to\infty}(T_{1})^{n}c\leq\lim_{n\to\infty}(T_{2})^{n}c=c_{2}^{*}$, which proves the claim once $T_{1}c\leq T_{2}c$ is established.
[…]
$u^{\prime}(\xi_{2},z)>u^{\prime}(\xi_{1},z)$ $=\max\{\mathbbm{E}_{z}\hat{\beta}\hat{R}u^{\prime}(c(\hat{R}(w-\xi_{1})+\hat{Y}_{1},\hat{Z}),\hat{Z}),u^{\prime}(w,z)\}$
$\geq\max\{\mathbbm{E}_{z}\,\hat{\beta}\hat{R}u^{\prime}(c(\hat{R}(w-\xi_{2})+\hat{Y}_{2},\hat{Z}),\hat{Z}),u^{\prime}(w,z)\}=u^{\prime}(\xi_{2},z),$
which is a contradiction. This completes the proof. ∎
Proof of Proposition 2.4.
Recall that, for all $c\in\mathscr{C}$, $\xi(w,z):=Tc(w,z)$ solves
(26) $u^{\prime}(\xi(w,z),z)=\max\left\{\mathbbm{E}_{z}\hat{\beta}\hat{R}u^{\prime}(c(\hat{R}\,[w-\xi(w,z)]+\hat{Y},\hat{Z}),\hat{Z}),u^{\prime}(w,z)\right\}.$
For each $z\in\mathsf{Z}$ and $c\in\mathscr{C}$, define
(27) $\bar{w}_{c}(z):=u^{\prime}(\cdot,z)^{-1}\left[\mathbbm{E}_{z}\hat{\beta}\hat{R}u^{\prime}(c(\hat{Y},\hat{Z}),\hat{Z})\right]\quad\text{and}\quad\bar{w}(z):=\bar{w}_{c^{*}}(z).$
Let $w\leq\bar{w}_{c}(z)$. We claim that $\xi(w,z)=w$. Suppose to the contrary that $\xi(w,z)<w$. Then $u^{\prime}(\xi(w,z),z)>u^{\prime}(w,z)$. In view of (26), we have
$u^{\prime}(w,z)<\mathbbm{E}_{z}\hat{\beta}\hat{R}u^{\prime}(c(\hat{R}\left[w-\xi(w,z)\right]+\hat{Y},\hat{Z}),\hat{Z})\leq\mathbbm{E}_{z}\hat{\beta}\hat{R}u^{\prime}(c(\hat{Y},\hat{Z}),\hat{Z})=u^{\prime}(\bar{w}_{c}(z),z).$
From this we get $w>\bar{w}_{c}(z)$, which is a contradiction. Hence, $\xi(w,z)=w$.
On the other hand, if $\xi(w,z)=w$, then $u^{\prime}(\xi(w,z),z)=u^{\prime}(w,z)$. By (26), we have $u^{\prime}(w,z)\geq\mathbbm{E}_{z}\,\hat{\beta}\hat{R}u^{\prime}(c(\hat{Y},\hat{Z}),\hat{Z})=u^{\prime}(\bar{w}_{c}(z),z)$. Hence, $w\leq\bar{w}_{c}(z)$. The first claim is verified. The second claim follows immediately from the first claim and the fact that $c^{*}$ is the unique fixed point of $T$ in $\mathscr{C}$. ∎
Appendix C Proofs of Theorems 3.13.2
To establish our key results, we extend the operator $F_{i}$ defined in (15) as follows. With slight abuse of notation, in what follows, for each $i\in\{1,\dots,N\}$ and $j\in\{1,\dots,M\}$, we define $F_{i}:[1,\infty]^{M}\to[1,\infty]^{M}$ by
(28) $(F_{i}x)(\tilde{z}_{j}):=\begin{cases}\infty,&\text{if\, $\sum_{\ell=1}^{i-1}\bar{p}_{i\ell}>0$},\\ \left(1+(G_{i}x)(\tilde{z}_{j})^{1/\gamma_{i}}\right)^{\gamma_{i}},&\text{if\, $\sum_{\ell=1}^{i-1}\bar{p}_{i\ell}=0$}.\end{cases}$
Moreover, with slight abuse of notation, we define
$\gamma_{i}=\gamma(\bar{z}_{i})=\gamma(z_{ij})\quad\text{for all }\,z_{ij}=(\bar{z}_{i},\tilde{z}_{j})\in\mathsf{Z}.$
Lemma C.1.
Fix $i\in\{1,\dots,N\}$. If $c\in\mathscr{C}$ and for all $j\in\{1,\dots,M\}$, we have
$\limsup_{w\to\infty}\frac{c(w,z_{ij})}{w}\leq x(\tilde{z}_{j})^{-1/\gamma_{i}},$
where $x(\tilde{z}_{j})\in[1,\infty]$, then for all $j\in\{1,\dots,M\}$, we have
(29) $\limsup_{w\to\infty}\frac{Tc(w,z_{ij})}{w}\leq(F_{i}x)(\tilde{z}_{j})^{-1/\gamma_{i}}.$
Proof.
Let $\alpha(z_{ij})=\limsup_{w\to\infty}Tc(w,z_{ij})/w$. By definition, we can take an increasing sequence $\{w_{n}\}$ such that $\alpha(z_{ij})=\lim_{n\to\infty}Tc(w_{n},z_{ij})/w_{n}$. Define $\alpha_{n}(z_{ij}):=Tc(w_{n},z_{ij})/w_{n}\in(0,1]$ and
(30) $\lambda_{n}(z_{ij})=c\big(\hat{R}(1-\alpha_{n}(z_{ij}))w_{n}+\hat{Y},\hat{Z}\big)\big/w_{n}>0.$
We next show that
(31) $\limsup_{n\to\infty}\lambda_{n}(z_{ij})\leq x(\hat{Z})^{-1/\gamma(\hat{Z})}\hat{R}(1-\alpha(z_{ij})).$
Note that
$\lambda_{n}(z_{ij})=\frac{c\big(\hat{R}(1-\alpha_{n}(z_{ij}))w_{n}+\hat{Y},\hat{Z}\big)}{\hat{R}(1-\alpha_{n}(z_{ij}))w_{n}+\hat{Y}}\left(\hat{R}(1-\alpha_{n}(z_{ij}))+\frac{\hat{Y}}{w_{n}}\right)$
If $\alpha(z_{ij})<1$ and $\hat{R}>0$, then since $\hat{R}(1-\alpha_{n}(z_{ij}))w_{n}\to\hat{R}(1-\alpha(z_{ij}))\cdot\infty=\infty$, by assumption we have
$\limsup_{n\to\infty}\lambda_{n}(z_{ij})\leq\limsup_{w\to\infty}\frac{c(w,\hat{Z})}{w}\hat{R}(1-\alpha(z_{ij}))\leq x(\hat{Z})^{-1/\gamma(\hat{Z})}\hat{R}(1-\alpha(z_{ij})),$
which is equation (31). If $\alpha(z_{ij})=1$ or $\hat{R}=0$, since $c(w,z)\leq w$, we have
$\lambda_{n}(z_{ij})\leq\hat{R}(1-\alpha_{n}(z_{ij}))+\hat{Y}/w_{n}\to\hat{R}(1-\alpha(z_{ij}))=0,$
so again equation (31) holds.
Since $Tc(w_{n},z_{ij})=\alpha_{n}(z_{ij})w_{n}$ solves the Euler equation, we have
$\alpha_{n}(z_{ij})^{-\gamma_{i}}$ $=\max\left\{\mathbbm{E}_{z_{ij}}\hat{\beta}\hat{R}\frac{c\big(\hat{R}(w_{n}-\alpha_{n}(z_{ij})w_{n})+\hat{Y},\hat{Z}\big)^{-\gamma(\hat{Z})}}{w_{n}^{-\gamma_{i}}},\;1\right\}$
$=\max\left\{\mathbbm{E}_{z_{ij}}\hat{\beta}\hat{R}\lambda_{n}(z_{ij})^{-\gamma(\hat{Z})}w_{n}^{\gamma_{i}-\gamma(\hat{Z})},\;1\right\}.$
Therefore,
(32) $\alpha_{n}(z_{ij})^{-\gamma_{i}}\geq\mathbbm{E}_{z_{ij}}\hat{\beta}\hat{R}\lambda_{n}(z_{ij})^{-\gamma(\hat{Z})}w_{n}^{\gamma_{i}-\gamma(\hat{Z})}$.
Case 1. If $\sum_{\ell=1}^{i-1}\bar{p}_{i\ell}>0$, then there exists at least one $h<i$ such that $\bar{p}_{ih}>0$. By definition, this implies that $\gamma_{i}-\gamma_{h}>0$ with positive probability. Letting $n\to\infty$ in (32) and applying Fatou’s lemma, we have
$\alpha(z_{ij})^{-\gamma_{i}}$ $=\lim_{n\to\infty}\alpha_{n}(z_{ij})^{-\gamma_{i}}\geq\liminf_{n\to\infty}\mathbbm{E}_{z_{ij}}\hat{\beta}\hat{R}\lambda_{n}(z_{ij})^{-\gamma(\hat{Z})}w_{n}^{\gamma_{i}-\gamma(\hat{Z})}$
[…]
$\alpha(z_{ij})^{-\gamma_{i}}$ $\geq\bar{p}_{ih}\tilde{p}_{jk}\mathbbm{E}_{z_{ij},z_{hk}}\left(\hat{\beta}\hat{R}\big[x(z_{hk})^{-1/\gamma_{h}}\hat{R}(1-\alpha(z_{ij}))\big]^{-\gamma_{h}}\lim_{n\to\infty}w_{n}^{\gamma_{i}-\gamma_{h}}\right)=\infty,$
where the expectation $\mathbbm{E}_{z_{ij},z_{hk}}$ is taken with respect to the innovation $\hat{\varepsilon}$, and the last equality holds because, in this case, $\mathbbm{P}_{(z,\hat{z})}(\hat{\beta}\hat{R}>0)>0$ for all $(z,\hat{z})\in\mathsf{Z}^{2}$ by assumption. In addition, since by definition $(F_{i}x)(\tilde{z}_{j})\equiv\infty$ in this case, we have
$\limsup_{w\to\infty}\frac{Tc(w,z_{ij})}{w}=\alpha(z_{ij})\leq 0=(F_{i}x)(\tilde{z}_{j})^{-1/\gamma_{i}}.$
Case 2. If $\sum_{\ell=1}^{i-1}\bar{p}_{i\ell}=0$ and $\bar{p}_{ii}>0$, then, considering only the transition of state $\bar{z}_{i}$ to itself next period, (32) implies that
$\alpha_{n}(z_{ij})^{-\gamma_{i}}\geq\bar{p}_{ii}\sum_{k=1}^{M}\tilde{p}_{jk}\mathbbm{E}_{z_{ij},z_{ik}}\left[\hat{\beta}\hat{R}\lambda_{n}(z_{ij})^{-\gamma_{i}}\right].$
Letting $n\to\infty$ and applying Fatou’s lemma, we get
$\alpha(z_{ij})^{-\gamma_{i}}$ $=\lim_{n\to\infty}\alpha_{n}(z_{ij})^{-\gamma_{i}}\geq\liminf_{n\to\infty}\bar{p}_{ii}\sum_{k=1}^{M}\tilde{p}_{jk}\mathbbm{E}_{z_{ij},z_{ik}}\hat{\beta}\hat{R}\lambda_{n}(z_{ij})^{-\gamma_{i}}$
$\geq\bar{p}_{ii}\sum_{k=1}^{M}\tilde{p}_{jk}\mathbbm{E}_{z_{ij},z_{ik}}\hat{\beta}\hat{R}\liminf_{n\to\infty}\lambda_{n}(z_{ij})^{-\gamma_{i}}$
$=\bar{p}_{ii}\sum_{k=1}^{M}\tilde{p}_{jk}\mathbbm{E}_{z_{ij},z_{ik}}\hat{\beta}\hat{R}\big(\limsup_{n\to\infty}\lambda_{n}(z_{ij})\big)^{-\gamma_{i}}$
$\geq\bar{p}_{ii}\sum_{k=1}^{M}\tilde{p}_{jk}\mathbbm{E}_{z_{ij},z_{ik}}\hat{\beta}\hat{R}\big[x(z_{ik})^{-1/\gamma_{i}}\hat{R}(1-\alpha(z_{ij}))\big]^{-\gamma_{i}}.$
Solving the above inequality, we have
$\limsup_{w\to\infty}\frac{Tc(w,z_{ij})}{w}$ $=\alpha(z_{ij})\leq\frac{1}{1+\left(\bar{p}_{ii}\sum_{k=1}^{M}\tilde{p}_{jk}\mathbbm{E}_{z_{ij},z_{ik}}\hat{\beta}\hat{R}^{1-\gamma_{i}}x(z_{ik})\right)^{1/\gamma_{i}}}$
$=(F_{i}x)(\tilde{z}_{j})^{-1/\gamma_{i}}.$
[…]
Lemma C.2.
Let $G$ be a $M\times M$ nonnegative matrix. Define $F:\mathbbm{R}_{+}^{M}\to\mathbbm{R}_{+}^{M}$ by $Fx=\varphi(Gx)$, where $\varphi(t)=(1+t^{1/\gamma})^{\gamma}$ and $\gamma>0$ is a constant. Then $F$ has a fixed point $x^{*}\in\mathbbm{R}_{+}^{M}$ if and only if $r(G)<1$, and the fixed point is unique. Take any $x_{0}\in\mathbbm{R}_{+}^{M}$ and define the sequence $\{x_{n}\}_{n=1}^{\infty}\subset\mathbbm{R}_{+}^{M}$ by $x_{n}=Fx_{n-1}$ for all $n\in\mathbb{N}$. Then we have the following:
  1. (1)
    If $r(G)<1$, then $\{x_{n}\}_{n=1}^{\infty}$ converges to $x^{*}$.
  2. (2)
    If $r(G)\geq 1$ and $G$ is irreducible, then $\{x_{n}\}_{n=1}^{\infty}$ converges to $x^{*}=\infty$.
Proof.
This follows from Lemmas 15–16 of Ma and Toda (2021). ∎
Proposition C.1.
For given $i\in\{1,\dots,N\}$, if $\sum_{\ell=1}^{i-1}\bar{p}_{i\ell}=0$ and $\bar{p}_{ii}>0$, then $F_{i}$ defined in (28) has a fixed point $x_{i}^{*}\in\mathbbm{R}_{+}^{M}$ if and only if $r(G_{i})<1$, and the fixed point is unique. Take any $x_{0}\in\mathbbm{R}_{+}^{M}$, and define the sequence $\{x_{n}\}_{n=1}^{\infty}\subset\mathbbm{R}_{+}^{M}$ by $x_{n}=F_{i}x_{n-1}$ for all $n\in\mathbb{N}$. Then we have the following:
  1. (1)
    If $r(G_{i})<1$, then $\{x_{n}\}_{n=1}^{\infty}$ converges to $x_{i}^{*}$.
  2. (2)
    If $r(G_{i})\geq 1$ and $G_{i}$ is irreducible, then $\{x_{n}\}_{n=1}^{\infty}$ converges to $x_{i}^{*}=\infty$.
Proof.
This follows immediately from Lemma C.2 by letting $F=F_{i}$ and $G=G_{i}$. ∎
Proof of Theorem 3.1.
Define the sequence $\{c_{n}\}\subset\mathscr{C}$ by $c_{0}(w,z)=w$ and $c_{n}(w,z)=Tc_{n-1}(w,z)$ for all $n\geq 1$. Since $Tc(w,z)\leq w$ for any $c\in\mathscr{C}$, we have $c_{1}(w,z)=Tc_{0}(w,z)\leq w=c_{0}(w,z)$. Since $T:\mathscr{C}\to\mathscr{C}$ is monotone, by induction $0\leq c_{n}(w,z)\leq c_{n-1}(w,z)\leq w$ for all $n\geq 1$ $S(w,z)\in\SS, and$ $c^{*}(w,z)=\lim_{n\to\infty}c_{n}(w,z)$ exists. This $c^{*}$ is the unique fixed point of $T$ and also the unique optimal policy in $\mathscr{C}$.
Fix $i\in\{1,\dots,N\}$. Define the sequence $\{x_{n}\}\subset\mathbbm{R}_{+}^{M}$ by
$x_{0}(\tilde{z}_{j})=1\quad\text{and}\quad x_{n}(\tilde{z}_{j})=(F_{i}x_{n-1})(\tilde{z}_{j}),\qquad j=1,\dots,M.$
By definition, $c_{0}(w,z_{ij})/w=1=x_{0}(\tilde{z}_{j})^{-1/\gamma_{i}}$. Hence, for all $j\in\{1,\dots,M\}$,
$\limsup_{w\to\infty}\frac{c_{0}(w,z_{ij})}{w}\leq x_{0}(\tilde{z}_{j})^{-1/\gamma_{i}}.$
Since $c_{n}(w,z_{ij})\downarrow c^{*}(w,z_{ij})$ pointwise, a repeated application of Lemma C.1 yields
(33) $0\leq\liminf_{w\to\infty}\frac{c^{*}(w,z_{ij})}{w}\leq\limsup_{w\to\infty}\frac{c^{*}(w,z_{ij})}{w}\leq\limsup_{w\to\infty}\frac{c_{n}(w,z_{ij})}{w}\leq x_{n}(\tilde{z}_{j})^{-1/\gamma_{i}}.$
Case 1. If $\sum_{j=1}^{i-1}\bar{p}_{ij}>0$, then Lemma C.1 implies that $x_{n}(\tilde{z}_{j})=\infty$ for all $n\geq 1$ and $j\in\{1,\dots,M\}$. Then (12) follows from (33).
Case 2. If $\bar{p}_{ii}>0$, $G_{i}$ is irreducible, and $r(G_{i})\geq 1$, then by Proposition C.1, we have $x_{n}(\tilde{z}_{j})\to\infty$ as $n\to\infty$ for all $j\in\{1,\dots,M\}$. Again, (12) follows from (33). ∎
Proof of Theorem 3.2.
If each entry of $\bar{P}$ is positive, then condition (1) of Theorem 3.1 holds. If in addition $G_{1}$ is irreducible and $r(G_{1})\geq 1$, then condition (2) of Theorem 3.1 holds. In both cases, so the conclusion follows trivially. ∎
Appendix D Proofs of Theorems 3.33.5
To prove Theorem 3.3, we need the following lemmas.
Lemma D.1.
Fix $i\in\{1,\dots,N\}$. Consider $F_{i}$ defined in (28) with $\bar{p}_{ii}=1$. Suppose $r(G_{i})<1$ and let $x_{i}^{*}\in\mathbbm{R}_{+}^{M}$ be the unique fixed point of $F_{i}$. If $c\in\mathscr{C}$ and for all $w>0$ and $j\in\{1,\dots,M\}$, we have
$\frac{c(w,z_{ij})}{w}\geq x_{i}^{*}(\tilde{z}_{j})^{-1/\gamma_{i}},$
then for all $w>0$ and $j\in\{1,\dots,M\}$, we have
$\frac{Tc(w,z_{ij})}{w}\geq x_{i}^{*}(\tilde{z}_{j})^{-1/\gamma_{i}}.$
Proof.
Suppose $Tc(w,z_{ij})/w<x_{i}^{*}(\tilde{z}_{j})^{-1/\gamma_{i}}$ for some $w$ and $z_{ij}$. Let $\xi:=Tc(w,z_{ij})$. Then $x_{i}^{*}(\tilde{z}_{j})^{-1/\gamma_{i}}w>\xi$. Since in addition $x_{i}^{*}(\tilde{z}_{j})^{-1/\gamma_{i}}\in(0,1]$ by Proposition C.1 (claim (1)), by the monotonicity of $u^{\prime}$, we have
$w^{-\gamma_{i}}$ $=u^{\prime}(w,z_{ij})\leq u^{\prime}(x_{i}^{*}(\tilde{z}_{j})^{-1/\gamma_{i}}w,z_{ij})<u^{\prime}(\xi,z_{ij})$
$=\max\left\{\mathbbm{E}_{z_{ij}}\hat{\beta}\hat{R}u^{\prime}(c(\hat{R}(w-\xi)+\hat{Y},\hat{Z}),\hat{Z}),\;u^{\prime}(w,z_{ij})\right\}$
$=\max\left\{\mathbbm{E}_{z_{ij}}\hat{\beta}\hat{R}c(\hat{R}(w-\xi)+\hat{Y},\hat{Z})^{-\gamma_{i}},\;w^{-\gamma_{i}}\right\}.$
[…]
$w^{-\gamma_{i}}<\mathbbm{E}_{z_{ij}}\hat{\beta}\hat{R}c(\hat{R}(w-\xi)+\hat{Y},\hat{Z})^{-\gamma_{i}}.$
Under the contradiction hypothesis, we have
$x_{i}^{*}(\tilde{z}_{j})w^{-\gamma_{i}}$ $<\xi^{-\gamma_{i}}=\mathbbm{E}_{z_{ij}}\hat{\beta}\hat{R}c(\hat{R}(w-\xi)+\hat{Y},\hat{Z})^{-\gamma_{i}}$
$\leq\mathbbm{E}_{z_{ij}}\hat{\beta}\hat{R}x_{i}^{*}(\hat{\tilde{Z}})(\hat{R}(w-\xi)+\hat{Y})^{-\gamma_{i}}$
$<\mathbbm{E}_{z_{ij}}\hat{\beta}\hat{R}^{1-\gamma_{i}}x_{i}^{*}(\hat{\tilde{Z}})\left(1-x_{i}^{*}(\tilde{z}_{j})^{-1/\gamma_{i}}\right)^{-\gamma_{i}}w^{-\gamma_{i}}.$
By simple algebra and using the fact that $\bar{p}_{ii}=1$, we obtain
$x_{i}^{*}(\tilde{z}_{j})$ $<\left(1+\left[\mathbbm{E}_{z_{ij}}\hat{\beta}\hat{R}^{1-\gamma_{i}}x_{i}^{*}(\hat{\tilde{Z}})\right]^{1/\gamma_{i}}\right)^{\gamma_{i}}$
$=\left(1+\left[\bar{p}_{ii}\sum_{k=1}^{M}\tilde{p}_{jk}\mathbbm{E}_{z_{ij},z_{ik}}\hat{\beta}\hat{R}^{1-\gamma_{i}}x_{i}^{*}(\tilde{z}_{k})\right]^{1/\gamma_{i}}\right)^{\gamma_{i}}=(F_{i}x_{i}^{*})(\tilde{z}_{j}).$
This contradicts the definition of $F_{i}$ and $x_{i}^{*}(\tilde{z}_{j})$. ∎
Proof of Theorem 3.3.
Same to the proof of Theorem 3.1, we define the sequence $\{c_{n}\}\subset\mathscr{C}$ by $c_{0}(w,z)=w$ and $c_{n}=Tc_{n-1}$ for all $n\geq 1$, which gives a decreasing sequence of functions $0\leq c_{n}\leq c_{n-1}\leq w$ for all $n$, and the optimal policy $c^{*}(w,z)=\lim_{n\to\infty}c_{n}(w,z)$ is well defined. Moreover, define the sequence $\{x_{n}\}\subset\mathbbm{R}_{+}^{M}$ by $x_{0}(\tilde{z}_{j})=1$ and $x_{n}(\tilde{z}_{j})=(F_{i}x_{n-1})(\tilde{z}_{j})$ for all $j\in\{1,\dots,M\}$. A repeated application of Lemma C.1 gives
(34) $\limsup_{w\to\infty}\frac{c^{*}(w,z_{ij})}{w}\leq x_{i}^{*}(\tilde{z}_{j})^{-1/\gamma_{i}}.$
Since $c_{0}(w,z_{ij})/w=1\geq x_{i}^{*}(\tilde{z}_{j})^{-1/\gamma_{i}}$ for all $w>0$ and $z_{ij}$, a repeated application of Lemma D.1 implies that $c_{n}(w,z_{ij})/w\geq x_{i}^{*}(\tilde{z}_{j})^{-1/\gamma_{i}}$ for all $w>0$ and $z_{ij}$. Since $c_{n}\to c^{*}$ pointwise, letting $n\to\infty$, dividing both sides by $w>0$, and letting $w\to\infty$, we obtain
(35) $\liminf_{w\to\infty}\frac{c^{*}(w,z_{ij})}{w}\geq x^{*}(\tilde{z}_{j})^{-1/\gamma_{i}}$
for all $j\in\{1,\dots,M\}$. Then the stated claim follows from (34) and (35). ∎
To prove Theorem 3.43.5, we need the following lemma.
Lemma D.2.
Fix $i\in\{1,\dots,N\}$. Suppose the following conditions hold:
  1. (1)
    $\sum_{j=1}^{i-1}\bar{p}_{ij}=0$.
  2. (2)
    There exists $m>0$ such that $R\geq m$ with probability one.
If $c\in\mathscr{C}$ and for all $z_{ij}\in\mathsf{Z}$, we have
$\lim_{w\to\infty}\frac{c(w,z_{ij})}{w}=x(\tilde{z}_{j})^{-1/\gamma_{i}},$
then for all $z_{ij}\in\mathsf{Z}$, we have
(36) $\lim_{w\to\infty}\frac{Tc(w,z_{ij})}{w}=(F_{i}x)(\tilde{z}_{j})^{-1/\gamma_{i}}.$
Proof.
Fix $i\in\{1,\dots,N\}$. If $\mathbbm{P}_{z_{ij}}(\hat{\beta}\hat{R}>0)=0$ for some $j$, then by the definition of $T$ and $G_{i}$, we have $Tc(w,z_{ij})\equiv w$ and $(G_{i}x)(\tilde{z}_{j})\equiv 0$. In this case, (36) holds trivially since $\lim_{w\to\infty}Tc(w,z_{ij})/w=1=(F_{i}x)(\tilde{z}_{j})^{-1/\gamma_{i}}$. In what follows, we consider $\mathbbm{P}_{z_{ij}}(\hat{\beta}\hat{R}>0)>0$ for all $j$.
In this case, if $x(\tilde{z}_{j})=\infty$ for some $j$, then (36) holds trivially by Lemma C.1. In what follows, we consider $x(\tilde{z}_{j})<\infty$ for all $j$.
Let $\alpha(z_{ij})$ be an accumulation point of $Tc(w,z_{ij})/w$ as $w\to\infty$. By Lemma C.1, we have $\alpha(z_{ij})\leq 1$. By definition, we can take an increasing sequence $\{w_{n}\}$ such that $\alpha(z_{ij})=\lim_{n\to\infty}Tc(w_{n},z_{ij})/w_{n}$. Define $\alpha_{n}(z_{ij})=Tc(w_{n},z_{ij})/w_{n}\in(0,1]$ and $\lambda_{n}(z_{ij})$ as in Lemma C.1. By the proof of Lemma C.1, we have
(37) $\lim_{n\to\infty}\lambda_{n}(z_{ij})=x(\hat{Z})^{-1/\gamma(\hat{Z})}\hat{R}(1-\alpha(z_{ij})).$
Since $Tc(w_{n},z_{ij})=\alpha_{n}w_{n}$ solves the Euler equation, we have
$\alpha_{n}(z_{ij})^{-\gamma_{i}}$ $=\max\left\{\mathbbm{E}_{z_{ij}}\hat{\beta}\hat{R}\frac{c(\hat{R}(w_{n}-\alpha_{n}(z_{ij})w_{n})+\hat{Y},\hat{Z})^{-\gamma(\hat{Z})}}{w^{-\gamma_{i}}},\;1\right\}$
(38) $=\max\left\{\mathbbm{E}_{z_{ij}}\hat{\beta}\hat{R}\bigg[\frac{c(\hat{R}(w_{n}-\alpha_{n}(z_{ij})w_{n})+\hat{Y},\hat{Z})}{w_{n}}\bigg]^{-\gamma(\hat{Z})}w_{n}^{\gamma_{i}-\gamma(\hat{Z})},\;1\right\}.$
Since $\hat{R}\geq m>0$ with probability one by assumption, (37) implies that
(39) $\liminf_{n\to\infty}\lambda_{n}(z_{ij})\geq x(\hat{Z})^{-1/\gamma(\hat{Z})}m(1-\alpha(z_{ij})).$
Case 1. $\bar{p}_{ii}=0$. If $\bar{p}_{ii}=0$, then $(F_{i}x)(\tilde{z}_{j})\equiv 1$. It suffices to show that $\alpha(z_{ij})=1$ for all $j$. Because in this case, we have $\alpha(z_{ij})=(F_{i}x)(\tilde{z}_{j})$, then the stated claim holds.
Suppose $\alpha(z_{ij})<1$ for some $j$, then $\lim_{n\to\infty}\alpha_{n}(z_{ij})<1$, implying that $\alpha_{n}(z_{ij})<1$ for all $n\geq M$ for some integer $M$. By (38), we have
(40) $\alpha_{n}(z_{ij})^{-\gamma_{i}}=\mathbbm{E}_{z_{ij}}\hat{\beta}\hat{R}\bigg[\frac{c(\hat{R}(w_{n}-\alpha_{n}(z_{ij})w_{n})+\hat{Y},\hat{Z})}{w_{n}}\bigg]^{-\gamma(\hat{Z})}w_{n}^{\gamma_{i}-\gamma(\hat{Z})}$.
Since we now work with $x(\tilde{z}_{j})<\infty$ for all $j$, as pointed out above. By the definition of $F_{i}$, we have $x(\tilde{z}_{j})\in[1,\infty)$ for all $j$. Hence, for any
$\underline{\lambda}\in\left(0,\min_{j}x(\tilde{z}_{j})^{-1/\gamma_{i}}m(1-\alpha(z_{ij}))\right)\quad\text{and}\quad\bar{\alpha}\in\left(\max_{j}\alpha(z_{ij}),1\right),$
there exists $N\in\mathbb{N}$ such that $\lambda_{n}\geq\underline{\lambda}$, $\alpha_{n}<\bar{\alpha}$, and $w_{n}\geq 1$ for all $n\geq N$ and $z_{ij}$. The integrand in the expectation in (38) is bounded above by $\max_{z,\hat{z}\in\mathsf{Z}}\hat{\beta}\hat{R}\underline{\lambda}^{-\gamma(\hat{z})}$$\infty$, which is integrable. Letting $n\to\infty$, applying the dominated convergence theorem, and using $\alpha(z_{ij})<1$, we obtain
$\alpha(z_{ij})^{-\gamma_{i}}=\lim_{n\to\infty}\mathbbm{E}_{z_{ij}}\hat{\beta}\hat{R}\lambda_{n}(z_{ij})^{-\gamma(\hat{Z})}w_{n}^{\gamma_{i}-\gamma(\hat{Z})}=\mathbbm{E}_{z_{ij}}\hat{\beta}\hat{R}\left[\lim_{n\to\infty}\lambda_{n}(z_{ij})\right]^{-\gamma(\hat{Z})}\lim_{n\to\infty}w_{n}^{\gamma_{i}-\gamma(\hat{Z})}$.
Since $\sum_{\ell=1}^{i}\bar{p}_{i\ell}=0$ implies $\sum_{h=i+1}^{N}\bar{p}_{ih}=1$, and because $\mathbbm{P}_{z_{ij}}(\hat{\beta}\hat{R}>0)>0$ in this case, and because
$\left[\lim_{n\to\infty}\lambda_{n}(z_{ij})\right]^{-\gamma(\hat{Z})}\leq\max_{\hat{Z}\in\mathsf{Z}}\underline{\lambda}^{-\gamma(\hat{Z})}$$\infty$,
together with the fact that $\lim_{n\to\infty}w_{n}^{\gamma_{i}-\gamma_{h}}=0$ for all $h>i$, we obtain the conclusion that $\alpha(z_{ij})^{-\gamma_{i}}=0$. This gives $\alpha(z_{ij})=\infty$, which is contradicted with the assumption that $\alpha(z_{ij})<1$. Hence, we must have $\alpha(z_{ij})=1$ for all $j$.
Case 2. $\bar{p}_{ii}>0$. If $\bar{p}_{ii}>0$, then by Lemma C.1 we obtain $\alpha(z_{ij})<1$ for all $j$. Indeed, in this case $\mathbbm{P}_{z_{ij}}(\hat{\beta}\hat{R}>0)>0$ for all $j$, which guarantees that each row of $G_{i}$ has at least one positive entry. Hence $(G_{i}x)(\tilde{z}_{j})>0$ for all $j$, and consequently $(F_{i}x)(\tilde{z}_{j})>1$ for all $j$. Hence (40) holds. This implies that
$\alpha(z_{ij})^{-\gamma_{i}}$ $=\bar{p}_{ii}\sum_{k=1}^{M}\tilde{p}_{jk}\mathbbm{E}_{z_{ij},z_{ik}}\hat{\beta}\hat{R}\left[\lim_{n\to\infty}\lambda_{n}(z_{ij})\right]^{-\gamma_{i}}$
$+\sum_{h=i+1}^{N}\bar{p}_{ih}\sum_{k=1}^{M}\tilde{p}_{jk}\mathbbm{E}_{z_{ij},z_{hk}}\left(\hat{\beta}\hat{R}\left[\lim_{n\to\infty}\lambda_{n}(z_{ij})\right]^{-\gamma_{h}}\lim_{n\to\infty}w_{n}^{\gamma_{i}-\gamma_{h}}\right)$
$=\bar{p}_{ii}\sum_{k=1}^{M}\tilde{p}_{jk}\mathbbm{E}_{z_{ij},z_{ik}}\hat{\beta}\hat{R}\left(x(\tilde{z}_{k})^{-1/\gamma_{i}}\hat{R}(1-\alpha(z_{ij}))\right)^{-\gamma_{i}},$
where the second equality follows by the same reasoning as in case 1. Solving the above equality, we obtain
$\lim_{w\to\infty}\frac{Tc(w,z_{ij})}{w}$ $=\alpha(z_{ij})=\frac{1}{1+\left(\bar{p}_{ii}\sum_{k=1}^{M}\tilde{p}_{jk}\mathbbm{E}_{z_{ij},z_{ik}}\hat{\beta}\hat{R}^{1-\gamma_{i}}x(\tilde{z}_{k})\right)^{1/\gamma_{i}}}$
$=\frac{1}{1+(G_{i}x)(\tilde{z}_{j})^{1/\gamma_{i}}}=(F_{i}x)(\tilde{z}_{j})^{-1/\gamma_{i}}.$
The proof is now complete. ∎
Proof of Theorem 3.4.
Same as before, let $c_{0}(w,z)\equiv w$ and $c_{n}=T^{n}c_{0}$. A repeated application of Lemma C.1 implies that, for all $z_{ij}\in\mathsf{Z}$, we have
$\limsup_{w\to\infty}\frac{c^{*}(w,z_{ij})}{w}\leq(F_{i}^{n}x)(\tilde{z}_{j}).$
A repeated application of Lemma D.2 implies that
$\liminf_{w\to\infty}\frac{c^{*}(w,z_{ij})}{w}=(F_{i}^{n}x)(\tilde{z}_{j}).$
Since in addition $\bar{p}_{ii}=0$, we have $F_{i}^{n}(\tilde{z}_{j})\equiv 1$ for all $n\geq 1$. The stated claim follows immediately from the above inequalities. ∎
Proof of Proposition 3.2.
Fix $z_{ij}\in\mathsf{Z}$ and denote $z=z_{ij}$. By Proposition 2.4, we know that $c^{*}(w,z)=w$ if and only if $w\leq\bar{w}(z)$. In particular, $c^{*}(w,z)<w$ for all $w>\bar{w}(z)$. Moreover, since $c^{*}\in\mathscr{C}$, there exists some $M<\infty$ such that
$\bar{w}(z)\leq u^{\prime}(\cdot,z)^{-1}\left[\mathbbm{E}_{z}\hat{\beta}\hat{R}\left(u^{\prime}(\hat{Y},\hat{Z})+M\right)\right]<\infty,$
where the first inequality is due to the definition of $\bar{w}(z)$ and the monotonicity of $u^{\prime}(\cdot,z)$, and the second inequality follows from Assumption 2.2. Hence, there exist some $w\in(\bar{w}(z),\infty)$ and $\varepsilon>0$ such that
$c^{*}(w,z)<w(1-\varepsilon)\quad\Longrightarrow\quad c^{*}(w,z)/w<1-\varepsilon.$
However, by Theorem 3.4, there exists some sufficiently large $w^{\prime}>w$ such that
$c^{*}(w^{\prime},z)/w^{\prime}\geq 1-\varepsilon.$
As a result, we have
$c^{*}(w,z)/w<c^{*}(w^{\prime},z)/w^{\prime}$
for some $w^{\prime}>w$. Hence $c^{*}(w,z)$ is not concave in $w$. The proof is complete. ∎
To prove Theorem 3.5, we extend the utility function to
$u(c,z_{ij})=\begin{cases}\psi(\bar{z}_{i})\frac{c^{1-\gamma(\bar{z}_{i})}}{1-\gamma(\bar{z}_{i})},&\text{if }\gamma(\bar{z}_{i})>0\text{ and }\gamma(\bar{z}_{i})\neq 1,\\ \psi(\bar{z}_{i})\log c,&\text{if }\gamma(\bar{z}_{i})=1,\end{cases}$
where $\psi_{i}:=\psi(\bar{z}_{i})>0$ denotes the weight on $\bar{z}_{i}$. We denote
$\psi:=(\psi_{1},\dots,\psi_{N})=(\psi(\bar{z}_{i}),\dots,\psi(\bar{z}_{N})).$
Furthermore, we introduce a new operator. Fix $i\in\{i,\dots,N\}$, we define the $M\times M$ matrix $D_{i}$ with entries
$d_{jk}=\tilde{p}_{jk}\sum_{h=1}^{N}\bar{p}_{ih}\psi_{h}\mathbbm{E}_{z_{ij},z_{hk}}\beta(z_{ij},z_{hk},\hat{\varepsilon})R(z_{ij},z_{hk},\hat{\varepsilon})^{1-\gamma_{i}},$
where, as before, the expectation is taken with respect to $\hat{\varepsilon}$. We then set
$U_{i}=\tilde{P}\circ D_{i}$
where $\tilde{P}\circ D_{i}$ as the Hadamard product of $\tilde{P}$ and $D_{i}$. For $y\in\mathbbm{R}_{+}^{M}$, we define
(41) $(J_{i}y)(\tilde{z}_{j}):=\left(1+(U_{i}y)(\tilde{z}_{j})^{1/\gamma_{i}}\right)^{\gamma_{i}}.$
Denote by $c^{*}(w,z;\psi,\gamma)$ the optimal consumption at $(w,z)$ when the weight vector is $\psi$ and $\gamma_{i}\equiv\gamma$ for all $i$. As usual, denote by $c^{*}(w,z)$ the optimal consumption at $(w,z)$ when $\psi_{i}\equiv 1$ for all $i$.
Lemma D.3.
Fix $i\in\{1,\dots,N\}$. Suppose $\sum_{\ell=1}^{i-1}\bar{p}_{i\ell}=0$ and $\gamma\equiv\gamma_{i}$. If $r(U_{i})<1$, then $J_{i}:\mathbbm{R}_{+}^{M}\to\mathbbm{R}_{+}^{M}$ and $J_{i}$ has a unique fixed point $y_{i}^{*}\in\mathbbm{R}_{+}^{M}$. Moreover, for all $j\in\{1,\dots,M\}$, we have
(42) $\lim_{w\to\infty}\frac{c^{*}(w,z_{ij};\psi,\gamma_{i})}{w}=y_{i}^{*}(\tilde{z}_{j})^{-1/\gamma_{i}}.$
Proof.
The first claim follows immediately from Lemma C.2, which implies that $J_{i}$ has a unique fixed point in $y_{i}^{*}\in\mathbbm{R}_{+}^{M}$. The proof of the second claim is similar to the proof of Theorem 3.3 and thus omitted. ∎
Let $T_{0}$ be the time iteration operator for the special case $\gamma_{i}\equiv\gamma$.
Lemma D.4.
Fix $i\in\{1,\dots,N\}$. Suppose $\sum_{\ell=1}^{i-1}\bar{p}_{i\ell}=0$ and there exists $b>0$ such that $Y(z,\hat{z},\hat{\varepsilon})\geq b$ almost surely conditional on (z,z^)($\hat{z}$). Choose $\psi$ such that $b^{-\gamma_{h}}/\psi_{h}\leq b^{-\gamma_{i}}/\psi_{i}$ for all $h\geq i$. Then $c^{*}(w,z_{ij};\psi,\gamma_{i})\leq c^{*}(w,z_{ij})$ for all $w\geq b$ and $z_{ij}\in\mathsf{Z}$.
Proof.
Consider the initial candidate $c(w,z)\equiv w$. Since
$(T_{0}^{n}c)(w,z_{ij})\to c^{*}(w,z_{ij};\psi,\gamma_{i})\quad\text{and}\quad(T^{n}c)(w,z_{ij})\to c^{*}(w,z_{ij}),$
for all $(w,z_{ij})$, to prove the stated claim, it suffices to show that
$(T_{0}^{n}c)(w,z_{ij})\leq(T^{n}c)(w,z_{ij})\quad\text{for all $n\in\mathbbm{N}$, $w\geq b$, and $z_{ij}\in\mathsf{Z}$}.$
Denote $c_{n}^{0}:=T_{0}^{n}c$ and $c_{n}:=T^{n}c$. The claim obviously holds for $n=0$. Suppose it holds for arbitrary $n$. Then $c_{n}^{0}(w,z_{ij})\leq c_{n}(w,z_{ij})$. It remains to verify that $c_{n+1}^{0}(w,z_{ij})\leq c_{n+1}(w,z_{ij})$ for all $w\geq b$ and $z_{ij}\in\mathsf{Z}$. Suppose this is not true. Then there exists $w\geq b$ and $z_{ij}\in\mathsf{Z}$ such that $c_{n+1}^{0}(w,z_{ij})>c_{n+1}(w,z_{ij})$. Then we have $1\geq c_{n+1}^{0}(w,z_{ij})/w>c_{n+1}(w,z_{ij})/w.$ By the definition of $T$ and the induction argument, we have
$c_{n+1}(w,z_{ij})^{-\gamma_{i}}$ $=\mathbbm{E}_{z_{ij}}\hat{\beta}\hat{R}c_{n}(\hat{R}(w-c_{n+1}(w,z_{ij}))+\hat{Y},\hat{Z})^{-\hat{\gamma}}$
$\leq\mathbbm{E}_{z_{ij}}\hat{\beta}\hat{R}c_{n}(\hat{R}(w-c_{n+1}^{0}(w,z_{ij}))+\hat{Y},\hat{Z})^{-\hat{\gamma}}$
(43) $\leq\mathbbm{E}_{z_{ij}}\hat{\beta}\hat{R}c_{n}^{0}(\hat{R}(w-c_{n+1}^{0}(w,z_{ij}))+\hat{Y},\hat{Z})^{-\hat{\gamma}}$.
To proceed, we show that as long as $m$ is chosen sufficiently small, we have
(44) $c_{n}^{0}(w,z)\geq b\quad\text{for all }n\geq 0,w\geq b,\text{ and }z\in\mathsf{Z}.$
To see this, recall that based on the monotonicity of $T$, we have $c_{n}^{0}\leq c_{n-1}^{0}$ for all $n\geq 1$. Recall $\bar{w}_{c}$ and $\bar{w}$ defined in (27). To simplify notation, we denote $\bar{w}_{n}:=\bar{w}_{c_{n}^{0}}$. Then based on the monotonicity of $u$, we have $\bar{w}_{n}\geq\bar{w}$ for all $n\geq 0$. Moreover, by definition, $\bar{w}(z)>0$ for all $z\in\mathsf{Z}$. Hence, we can choose $b$ such that $b\in\left(0,\min_{z\in\mathsf{Z}}\bar{w}(z)\right)$. Then we have $c_{n}^{0}(b,z)=b$ by Proposition 2.4. Then, by the monotonicity of $c_{n}^{0}$, we have $c_{n}^{0}(w,z)\geq c_{n}^{0}(b,z)=b>0$ for all $w\geq b,z\in\mathsf{Z}$, and $n\geq 0$. This combined with (43) implies that
$c_{n+1}(w,z_{ij})^{-\gamma_{i}}$ $\leq\mathbbm{E}_{z_{ij}}\hat{\beta}\hat{R}b^{-\hat{\gamma}}\left[\frac{c_{n}^{0}(\hat{R}(w-c_{n+1}^{0}(w,z_{ij}))+\hat{Y},\hat{Z})}{b}\right]^{-\hat{\gamma}}$
$=\bar{p}_{ii}\sum_{k=1}^{M}\tilde{p}_{jk}\mathbbm{E}_{z_{ij},z_{ik}}\hat{\beta}\hat{R}b^{-\gamma_{i}}\left[\frac{c_{n}^{0}(\hat{R}(w-c_{n+1}^{0}(w,z_{ij}))+\hat{Y},z_{ik})}{b}\right]^{-\gamma_{i}}$
$+\sum_{h=i+1}^{N}\bar{p}_{ih}\sum_{k=1}^{M}\tilde{p}_{jk}\mathbbm{E}_{z_{ij},z_{hk}}\hat{\beta}\hat{R}b^{-\gamma_{h}}\left[\frac{c_{n}^{0}(\hat{R}(w-c_{n+1}^{0}(w,z_{ij}))+\hat{Y},z_{hk})}{b}\right]^{-\gamma_{h}}$
$\leq\bar{p}_{ii}\sum_{k=1}^{M}\tilde{p}_{jk}\mathbbm{E}_{z_{ij},z_{ik}}\hat{\beta}\hat{R}b^{-\gamma_{i}}\left[\frac{c_{n}^{0}(\hat{R}(w-c_{n+1}^{0}(w,z_{ij}))+\hat{Y},z_{ik})}{b}\right]^{-\gamma_{i}}$
$+\sum_{h=i+1}^{N}\bar{p}_{ih}\sum_{k=1}^{M}\tilde{p}_{jk}\mathbbm{E}_{z_{ij},z_{hk}}\hat{\beta}\hat{R}\frac{{\psi}_{h}b^{-\gamma_{i}}}{{\psi}_{i}}\left[\frac{c_{n}^{0}(\hat{R}(w-c_{n+1}^{0}(w,z_{ij}))+\hat{Y},z_{hk})}{b}\right]^{-\gamma_{i}}$
$\leq c_{n+1}^{0}(w,z_{ij})^{-\gamma_{i}},$
where first inequality follows from (43), and the second inequality follows from (44) and the fact that $b^{-\gamma_{h}}/\psi_{h}\leq b^{-\gamma_{i}}/\psi_{i}$. Hence $c_{n+1}(w,z_{ij})\geq c_{n+1}^{0}(w,z_{ij})$. This is a contradiction. The stated claim then follows by induction. ∎
Proof of Theorem 3.5.
By Proposition C.1, $F_{i}$ has a unique fixed point $x_{i}^{*}\in\mathbbm{R}_{+}^{M}$. By Lemma C.1, we have
(45) $\limsup_{w\to\infty}\frac{c^{*}(w,z_{ij})}{w}\leq x_{i}^{*}(\tilde{z}_{j})^{-1/\gamma_{i}}.$
Recall that, by assumption, $Y\geq b$ for some $b>0$ with probability one. We define
(46) $\psi_{h}=\begin{cases}1,&\text{if }h<i,\\ \alpha b^{\gamma_{i}-\gamma_{h}},&\text{if }h\geq i,\end{cases}$
where $\alpha$ can be arbitrarily small. By doing this, we can always choose $\psi$ such that the condition $r(U_{i})<1$ in Lemma D.3 holds. By Lemma D.3, $J_{i}$ has a unique fixed point $y_{i}^{*}\in\mathbbm{R}_{+}^{M}$.
Moreover, note that by the definition of $\psi_{h}$ in (46), we have $\psi_{i}=\alpha$ and $\psi_{h}=\psi_{i}b^{\gamma_{i}-\gamma_{h}}$ for $h>i$. Hence, $b^{-\gamma_{h}}/\psi_{h}=b^{-\gamma_{i}}/\psi_{i}$. By the parametric monotoniciy result from Lemma D.4, we have
(47) $c^{*}(w,z_{ij})\geq c^{*}(w,z_{ij};\psi,\gamma_{i}).$
Define the sequence $\{x_{n}\}\subset\mathbbm{R}_{+}^{M}$ by $x_{0}=y_{i}^{*}$ and $x_{n}=F_{i}x_{n-1}$. Then, by Lemma D.3,
$\lim_{w\to\infty}\frac{c^{*}(w,z_{ij};\psi,\gamma_{i})}{w}=x_{0}(\tilde{z}_{j})^{-1/\gamma_{i}}.$
Since by assumption $R\geq m$ for some $m>0$, Lemma D.2 applies. Applying $T^{n}$ to both sides of (47) and using Lemma D.2 repeatedly, we obtain
$\liminf_{w\to\infty}\frac{T^{n}c^{*}(w,z_{ij})}{w}\geq\lim_{w\to\infty}\frac{T^{n}c^{*}(w,z_{ij};\psi,\gamma_{i})}{w}=x_{n}(\tilde{z}_{j})^{-1/\gamma_{i}}.$
Letting $n\to\infty$, since $x_{n}(\tilde{z}_{j})\to x_{i}^{*}(\tilde{z}_{j})$ and $c^{*}=T^{n}c^{*}$, we obtain
(48) $\liminf_{w\to\infty}\frac{c^{*}(w,z_{ij})}{w}\geq x_{i}^{*}(\tilde{z}_{j})^{-1/\gamma_{i}}.$
Then the stated claim follows from (45) and (48). ∎
Appendix E Quantitative Algorithms
In this section, we present the quantitative algorithms used in Section 4. We first describe the endogenous grid method for solving the optimal consumption policy, followed by the algorithms for computing state-dependent impulse responses.
E.1. The Endogenous Grid Algorithm
The model is solved using a generalized endogenous grid method (Carroll, 2006), which avoids root-finding during operator iterations by fixing a grid for saving rather than wealth. Specifically, we construct a finite saving grid $\mathcal{S}_{G}:=\{s_{g}\}_{g=1}^{G}$, where $0=s_{1}<\cdots<s_{G}$, update a candidate consumption function $c$ on $\mathcal{S}_{G}\times\mathsf{Z}$, and then determine the corresponding wealth grid endogenously based on optimal consumption and saving.
For each $t\in\mathbbm{N}$, let $\mathcal{W}^{t}_{G}:=\left\{w^{t}_{g}(z)\right\}_{(g,z)=(1,1)}^{(G,Z)}$ denote the endogenous wealth grid at iteration $t$, where $w^{t}_{g}(z)$ is the wealth corresponding to saving $s_{g}$ and exogenous state $z$. Set $w^{0}_{g}(z)=s_{g}$ for all $g$ and $z$. Let $\mathscr{C}(\mathcal{W}^{t}_{G})$ denote the set of continuous piecewise linear functions $c:(0,\infty)\times\mathsf{Z}\to\mathbbm{R}$ such that for each $z\in\mathsf{Z}$: (1) $0<c(w,z)\leq w$ for all $w>0$; (2) $c(w,z)=w$ for $0<w\leq w^{t}_{1}(z)$; (3) $c(w,z)$ is linear in $w$ on each subinterval $w^{t}_{g}(z),w^{t}_{g+1}(z)]$ for g=1,,G1g=1,...,G-1; and (4) $c(w,z)$ is linearly extrapolated for $w>w^{t}_{G}(z)$.
Algorithm 1   The endogenous grid algorithm
  1. Step 1.
    (Initialization). Choose a convergence criterion $\varpi>0$, a saving grid $\left\{s_{g}\right\}$, and an initial policy $c_{0}(w,z)$.
  2. Step 2.
    (Updating). For each $t\in\mathbbm{N}$, given $c_{t-1}\in\mathscr{C}(\mathcal{W}^{t-1}_{G})$, update it as follows:
    1. (a)
      For each $(s_{g},z)\in\mathcal{S}_{G}\times\mathsf{Z}$, compute $\hat{w}:=\hat{R}s_{g}+\hat{Y}$ and define
[…]
    2. (b)
      Define the upated wealth grid $\mathcal{W}^{t}_{G}:=\left\{w^{t}_{g}(z)\right\}_{(g,z)=(1,1)}^{(G,Z)}$ and the optimal consumption on it by
      $w^{t}_{g}(z):=s_{g}+\tilde{c}_{t}(s_{g},z)\quad\text{and}\quad c_{t}(w^{t}_{g}(z),z):=\tilde{c}_{t}(s_{g},z).$
    3. (c)
      Extend $c_{t}$ to $c_{t}\in\mathscr{C}(\mathcal{W}^{t}_{G})$ by linear interpolation and extrapolation and set $c_{t}(w,z)=w$ for $w\leq w^{t}_{1}(z)$.
  3. Step 3.
    (Convergence). Repeat Step 2 until $\left\{c_{t}(w^{t}_{g}(z),z)\right\}_{(g,z)=(1,1)}^{(G,Z)}$ converge.
The endogenous grid algorithm for computing the consumption function is summarized by Algorithm 1. In our quantitative experiments, we use an exponential savings grid with minimum, median, and maximum values of 0, $50$, and $10^{5}$, respectively, and 1000 grid points. The method for constructing the exponential grid follows Appendix B of Ma and Toda (2022). Iterations are terminated at precision $\varpi=10^{-4}$.
E.2. The Generalized Impulse Response Functions
Suppose a shock to the exogenous state process $\{Z_{t}\}$ occurs at the beginning of period $t$. As noted earlier, we follow Koop et al. (1996) and compute impulse response functions (IRFs) as state-dependent random variables, conditional on different initial states $(w_{t-1},Z_{t-1})$. To proceed, we define
[…]
Algorithm 2  The generalized impulse response function
  1. Step 1.
    (Initialization). Choose initial values $(w_{t-1},Z_{t-1})$, the horizon $H$, and the number of Monte Carlo samples $K$. Set
    $(\check{w}^{k}_{t-1},\check{Z}^{k}_{t-1})=(w_{t-1}^{k},Z^{k}_{t-1})\equiv(w_{t-1},Z_{t-1})\quad\text{for $k=1,\dots,K$.}$
  2. Step 2.
    Randomly sample $(H+1)\times K$ values of income shocks $\{\varepsilon^{k}_{t+h}\}_{(h,k)=(0,1)}^{(H,K)}$.
  3. Step 3.
    (Baseline Economy). Sample $(H+1)\times K$ values of the exogenous state
    $\left\{Z^{k}_{t+h}\right\}_{(h,k)=(0,1)}^{(H,K)}\quad\text{where }\;Z^{k}_{t+h}\sim P(Z^{k}_{t+h-1},\cdot\,).$
  4. Step 4.
    (Impulse Shock Economy). Compute the period-$t$ exogenous state $\{\check{Z}^{k}_{t}\}_{k=1}^{K}$ after the shock. Sample $H\times K$ values of risk aversion
    $\left\{\check{Z}^{k}_{t+h}\right\}_{(h,k)=(1,1)}^{(H,K)}\quad\text{where }\;\check{Z}^{k}_{t+h}\sim P(\check{Z}^{k}_{t+h-1},\cdot\,).$
  5. Step 5.
    For h=0,,Hh=0,...,H and k=1,,Kk=1,...,K, compute the sequence of wealth
    $w_{t+h}^{k}=F(w_{t+h-1}^{k},Z_{t+h-1}^{k},Z_{t+h}^{k},\varepsilon_{t+h}^{k})$
    $\check{w}_{t+h}^{k}=F(\check{w}_{t+h-1}^{k},\check{Z}_{t+h-1}^{k},\check{Z}_{t+h}^{k},\varepsilon_{t+h}^{k}).$
  6. Step 6.
    For h=0,,Hh=0,...,H, compute the period-$(t+h)$ impulse response
    $\text{IRF}(t+h)=\frac{1}{K}\sum_{k=1}^{K}c^{*}(\check{w}_{t}^{k},\check{Z}_{t}^{k})-\frac{1}{K}\sum_{k=1}^{K}c^{*}(w_{t}^{k},Z_{t}^{k}).$
Algorithm 2 summarizes the computation of generalized IRFs, where we use the absolute consumption response for demonstration. The method can be easily adapted to compute percentage changes, as well as impulse responses for saving rates and consumption volatility.
Recall the transition matrices $\bar{P}$ for $\{Z_{t}\}$ and $\tilde{P}$ for $\{\tilde{Z}_{t}\}$. In the quantitative experiments of Section 4.3, we have
$Z_{t}=(\bar{Z}_{t},\tilde{Z}_{t})=(\gamma_{t},\eta_{t}),\quad\varepsilon_{t}=(\nu_{t},\varepsilon_{t}^{\eta})\quad\text{and}\quad P(Z_{t-1},Z_{t})=\bar{P}(\gamma_{t-1},\gamma_{t})\tilde{P}(\eta_{t-1},\eta_{t}).$
A one-unit shock is applied to $\gamma_{t}$ at the beginning of period $t$. We discretize $\{\gamma_{t}\}$ into a 121-state Markov chain and $\{\eta_{t}\}$ into a 5-state Markov chain using the method of Tauchen (1986). To compute the period-$t$ exogenous state in the impulse shock economy (Step 4 of Algorithm 2), we search the $\gamma_{t}$ states and identify the smallest $\gamma_{t}$ that is greater than or equal to $\mu_{\gamma}(1-\rho_{\gamma})+\rho_{\gamma}\gamma_{t-1}+1$. Given the fine discretization of $\{\gamma_{t}\}$, the resulting error is negligible.